### v1

Abstract
This paper serves as a comprehensive system description of version 2.0 of the Marabou framework for formal analysis of neural networks. We discuss the tool’s architectural design and highlight the major features and components introduced since its initial release.
1 Introduction
With the increasing pervasiveness of deep neural networks (DNNs), the formal analysis of DNNs has become a burgeoning research field within the formal methods community. Multiple DNN reasoners have been proposed in the past few years, including $\alpha$-$\beta$-CROWN [76, 66, 73], ERAN [55, 56, 57], Marabou [40], MN-BaB [18], NNV [61, 44], nnenum [4], VeriNet [30, 29], and many others.
We focus here on the Marabou [40] tool, which has been used by the research community in a wide range of formal DNN reasoning applications (e.g., [26, 12, 46, 19, 59, 72, 9, 20, 64, 74, 32, 43], inter alia). Initially, Marabou was introduced as a from-scratch re-implementation of the Reluplex [39] decision procedure, with a native linear programming engine and limited support for DNN-level reasoning. Over the years, fundamental changes have been made to the tool, not only from an algorithmic perspective but also to its engineering and implementation.
This paper introduces version 2.0 of Marabou. Compared to its predecessor, Marabou 2.0: (i) employs a new build/test system; (ii) has an optimized core system architecture; (iii) runs an improved decision procedure and abstract interpretation techniques; (iv) handles a wider range of activation functions; (v) supports proof production; (vi) supports additional input formats; and (vii) contains a more powerful Python API. Due to these changes, the original system description [40] no longer gives an accurate account of the tool’s current capabilities. Our goal in this paper is to close this gap and provide a comprehensive description of the current Marabou system. We highlight the major features introduced since the initial version, describe a few of its many recent uses, and report on its performance, as demonstrated by the VNN-COMP’23 results and additional runtime comparisons against an early version of Marabou.
2 Architecture and Core Components
C++ APICLIPython APICEGAR SolverMulti-thread ManagerAbstract InterpretationsSimplex EngineSMT Solver(MI)LP InterfaceLinear Solving EngineSMT CoreContext-dependentData StructuresNetwork-level ReasonerPreprocessorProof ModuleBack EndEngineNon-linear Constraints(In)equationsPiecewise-linear ConstraintsBoundsInput QueryFront End
Figure 1: High-level overview of Marabou 2.0’s system architecture.
In this section, we discuss the core components of Marabou 2.0. An overview of its system architecture is given in Figure 1. At a high level, Marabou performs satisfiability checking on a set of linear and non-linear constraints, supplied through one of the front-end interfaces. The constraints typically represent a verification query over a neural network and are stored in an InputQuery object. We distinguish variable bounds from other linear constraints, and piecewise-linear constraints (which can be reduced to linear constraints via case analysis) from more general, non-linear constraints.
Variables are represented as consecutive indices starting from 0. (In)equations are represented as Equation objects. Piecewise-linear constraints are represented by objects of classes that inherit from the PiecewiseLinearConstraint abstract class. The abstract class defines the key interface methods that are implemented in each sub-class. This way, all piecewise-linear constraints are handled uniformly in the back end. Similarly, each other type of non-linear constraint is implemented as a sub-class of the new NonlinearConstraint abstract class. Initially, Marabou only supported the ReLU and Max constraints. In Marabou 2.0, over ten types of non-linear constraints (listed in Appendix  0.A.1) are supported.
2.1 Engine
The centerpiece of Marabou is called the Engine, which reasons about the satisfiability of the input query. The engine consists of several components: the Preprocessor, which performs rewrites and simplifications; the Network-level Reasoner, which maintains the network architecture and performs all analyses that require this knowledge; the SMT Solver, which houses complete decision procedures for sets of linear and piecewise-linear constraints; and the (MI)LP Interface, which manages interactions with external (MI)LP solvers for certain optional solving modes as explained below.
Two additional modules are built on top of the Engine. The Multi-thread Manager spawns multiple Engine instances to take advantage of multiple processors. The CEGAR Solver performs incremental linearization [70, 13] for non-linear constraints that cannot be precisely handled by the SMT Solver.
2.1.1 Preprocessor.
Every verification query first goes through multiple preprocessing passes, which normalize, simplify, and rewrite the query. One new normalizing pass introduces auxiliary variables and entailed linear constraints for each of the piecewise-linear constraints, so that case splits on the piecewise-linear constraints can be represented as bound updates and consequently do not require adding new equations. [^1] For example, for a piece-wise linear constraint $y=\max(x_{1},x_{2})$, we would introduce $c_{1}:y-x_{1}=a_{1}\land a_{1}\geq 0\land y-x_{2}=a_{2}\land a_{2}\geq 0$, where $a_{1}$ and $a_{2}$ are fresh variables. This way, case splits on this constraint can be represented as $c_{2}:a_{1}\leq 0$ and $c_{3}:a_{2}\leq 0$, respectively. This preprocessing pass preserves satisfiability because the original constraint is equisatisfiable to $c_{1}\land(c_{2}\lor c_{3})$. This accelerates the underlying Simplex engine, as explained in the SMT Solver section below. Another significant preprocessing pass involves iterative bound propagation over all constraints. In this process, piecewise linear constraints might collapse into linear constraints and be removed. This pass was present in Marabou 1.0, but could become a runtime bottleneck; whereas Marabou 2.0 employs a data structure optimization that leads to a $\sim$60x speed up. Finally, the preprocessor merges any variables discovered to be equal to each other and also eliminates any constant variables. This results in updates to the variable indices, and therefore a mapping from old indices to new ones needs to be maintained for retrieving satisfying assignments.
2.1.2 SMT Solver.
The SMT Solver module implements a sound and complete, lazy-DPLL(T)-based procedure for deciding the satisfiability of a set of linear and piecewise-linear constraints. It performs case analysis on the piecewise-linear constraints and, at each search state, employs a specialized procedure to iteratively search for an assignment satisfying both the linear and non-linear constraints.
Presently, the DeepSoI procedure [68] has replaced the Reluplex procedure [39, 40] as Marabou’s default procedure to run at each search state. The former provably converges to a satisfying assignment (if it exists) and empirically consistently outperforms the latter. DeepSoI extends the canonical sum-of-infeasibilities method in convex optimization [10], which determines the satisfiability of a set of linear constraints by minimizing a cost function that represents the total violation of the constraints by the current assignment. The constraints are satisfiable if and only if the optimal value is 0. Similarly, DeepSoI formulates a cost function that represents the total violation of the current piecewise-linear constraints and uses a convex solver to stochastically minimize the cost function with respect to the convex relaxation of the current constraints. In addition, DeepSoI also informs the branching heuristics of the SMT Core, which performs a case split on the piecewise-linear constraint with the largest impact (measured by the pseudocost metric [68]) on the cost function. The DeepSoI procedure is implemented for all supported piecewise-linear activation functions. The convex solver can be instantiated either with the native Simplex engine or with an external LP solver via the (MI)LP interface (detailed below). The latter can be more efficient but requires the use of external commercial solvers.
One optimization in Marabou 2.0’s Simplex engine is that once the tableau has been initialized, it avoids introducing any new equations — a costly operation that requires re-computing the tableau from scratch. This is achieved by implementing case-splitting and backtracking as updates on variable bounds (as mentioned above), which only requires minimal updates to the tableau state. By our measure, this optimization reduces the runtime of the Simplex engine by over 50%. Moreover, the memory footprint of the solver is also drastically decreased, as the SMT Core no longer needs to save the entire tableau state during case-splitting (to be restored during backtracking).
2.1.3 Network-level Reasoner.
Over the past few years, numerous papers (e.g.,  [65, 56, 77], inter alia) have proposed abstract interpretation techniques that rely on network-level reasoning (e.g., propagating the input bounds layer by layer to tighten output bounds). These analyses can be viewed as a stand-alone, incomplete DNN verification procedure, or as in-processing bound tightening passes for the SMT Solver. Marabou 2.0 features a brand new NetworkLevelReasoner class that supports this type of analysis. The class maintains the neural network topology as a directed acyclic graph, where each node is a Layer object. The Layer class records key information such as weights, source layers, and mappings between neuron indices and variable indices. Currently, seven different analyses are implemented: [i] 1. interval bound propagation [24]; 2. symbolic bound propagation [65]; 3. DeepPoly/CROWN analysis [56, 77]; 4. LP-based bound tightening [60]; 5. Forward-backward analysis [69]; 6. MILP-based bound tightening [60]; and 7. iterative propagation [67]. Analyses 2–7 are implemented in a parallelizable manner, and analyses 4–7 require calls to an external LP solver. By default, the DeepPoly/CROWN analysis is performed. The Network-level Reasoner tightly interleaves with the SMT Solver: the network-level reasoning is executed any time a new search state is reached (with the most up-to-date variable bounds), and the derived bound tightenings are immediately fed back to the search procedure.
It is noteworthy that the user does not have to explicitly provide the neural network topology to enable network-level reasoning. Instead, the network architecture is automatically inferred from the given set of linear and non-linear constraints, via the constructNetworkLevelReasoner method in the InputQuery class. The Network-level Reasoner is only initialized if such inference is successful. Apart from the abstract interpretation passes, the Network-level Reasoner can also evaluate concrete inputs. This is used to implement the LP-based bound tightening optimization introduced by the NNV tool [61].
2.1.4 (MI)LP Interface.
Marabou can now optionally be configured to invoke the Gurobi Optimizer [27], a state-of-the-art Mixed Integer Linear Programming (MILP) solver. The GurobiWrapper class contains methods to construct a MILP problem and invoke the solver. The MILPEncoder class is in charge of encoding the current set of linear and non-linear constraints as (MI)LP constraints. Piecewise-linear constraints can either be encoded precisely, or replaced with a convex relaxation, resulting in a linear program. For other non-linear constraints, only the latter option is available. The (MI)LP interface presently has three usages in the code base. Two have already been mentioned, i.e., in some of the abstract interpretation passes and optionally in the DeepSoI procedure. Additionally, when Marabou is compiled with Gurobi, a --milp mode is available, in which the Engine performs preprocessing and abstract interpretation passes, and then directly encodes the verification problem as a MILP problem to be solved by Gurobi. The mode is motivated by the observation that the performance of Gurobi and the SMT Solver can be complementary [68, 58].
2.1.5 Multi-thread Manager.
Parallelization is an important way to improve verification efficiency. Marabou supports two modes of parallelization, both managed by the new MultiThreadManager class: the split-and-conquer mode [67] and the portfolio mode. In the split-and-conquer mode, the original query is dynamically partitioned and re-partitioned into independent sub-queries, to be handled by idle workers. The partitioning strategy is implemented as a sub-class of the QueryDivider abstract class. Currently, two strategies are available: one partitions the intervals of the input variables; the other splits on piecewise linear constraints. By default, the former is used only when the input dimension is less than or equal to ten. In the portfolio mode, each worker solves the same query with a different random seed, which takes advantage of the stochastic nature of the DeepSoI procedure. Developing an interface to define richer kinds of portfolios is work in progress.
2.1.6 CEGAR Solver.
While the DNN verification community has by and large focused on piecewise-linear activation functions, other classes of non-linear connections exist and are commonly used for certain architectures [33, 63]. Apart from introducing support for non-linear constraints in the Preprocessor and the Network-level Reasoner, the latest Marabou version also incorporates a counter-example guided abstraction refinement (CEGAR) solving mode [70], based on incremental linearization [13] to enable more precise reasoning about non-linear constraints that are not piecewise linear. Currently, the CEGAR solver only supports Sigmoid and Tanh, but the module can be extended to handle other activation functions.
2.2 Context-Dependent Data-Structures
When performing a case split or backtracking to a previous search state, the SMT Core needs to save or restore information such as variable bounds and the phase status of each piecewise-linear constraint (e.g., is a ReLU currently active, inactive, or unfixed). To efficiently support these operations, Marabou 2.0 uses the notion of a context level (borrowed from the CVC4 SMT solver [6]), and stores the aforementioned information in context-dependent data structures. These data structures behave similarly to their standard counterparts, except that they are associated with a context level and automatically save and restore their state as the context increases or decreases. This major refactoring has greatly simplified the implementation of saving and restoring solver states and is an important milestone in an ongoing effort to integrate a full-blown Conflict-Driven Clause-Learning (CDCL) mechanism into Marabou.
2.3 Proof Module
A proof module has recently been introduced into Marabou, enabling it to optionally produce proof certificates after an unsatisfiable (UNSAT) [36] result. This is common practice in the SAT and SMT communities and is aimed at ensuring solver reliability. Marabou produces proof certificates based on a constructive variant of the Farkas lemma [62], which ensures the existence of a proof vector that witnesses the unsatisfiability of a linear program. Specifically, the proof vector corresponds to a linear equation that is violated by the variable bounds [36]. The full certificate of UNSAT is comprised of a proof tree, whose nodes represent the search states explored during the solving. Each node may contain a list of lemmas that are used as additional constraints in its descendent nodes; and each leaf node contains the proof vector for the unsatisfiability of the corresponding sub-query. The lemmas encapsulate some of the variable bounds, newly derived by the piecewiese-linear constraints of the query, and require their own witnesses (i.e., proof vectors). The BoundExplainer class is responsible for constructing all proof vectors, for updating them during execution, and for appending them to the node. The proof tree itself is implemented using the UnsatCertificateNode class.
When the solver is run in proof-production mode, the Proof module closely tracks the steps of the SMT Solver module and constructs the proof tree on the fly: new nodes are added to the tree whenever a case split is performed; and a new proof vector is generated whenever a lemma is learned or UNSAT is derived for a sub-query. If the Engine concludes that the entire query is UNSAT, a proof checker (implemented as an instance of the Checker class) will be triggered to certify the proof tree. It does so by traversing the tree and certifying the correctness of the lemmas and the unsatisfiability of the leaf nodes. A formally verified and precise proof-checker is currently under development [16]. Note that, currently, proof production mode is only compatible with a subset of the features supported by Marabou. Adding support for the remaining features (e.g., for the parallel solving mode) is an ongoing endeavor.
2.4 Front End
Marabou provides interfaces to prepare input queries and invoke the back-end solver in multiple ways. The Marabou executable can be run on the command line, taking in network/property/query files in supported formats. The Python and C++ APIs support this functionality as well, but also contain methods to add arbitrary linear and (supported) non-linear constraints. In addition, a layer on top of the Python API was added to Marabou 2.0 which allows users to define constraints in a more Pythonic manner, resulting in more succinct code. An example is given in Appendix. 0.A.2.
Typically, a query consists of the encoding of (one or several) neural networks and the encoding of a property on the network(s). To encode a neural network, the user has two options: 1) pass in a neural network file to be parsed by one of the neural network parsers; or 2) manually add constraints to encode the neural network. The main network format for Marabou 2.0 is now ONNX, towards which the neural network verification community is converging. The NNet format and the Tensorflow protobuf format are still supported but will likely be phased out in the long run. To encode the property on top of the neural network encoding, the user can 1) pass in a property file to be parsed by one of the property parsers; or 2) manually encode the property. Currently Marabou has two property parsers, one for a native property file format [40], and a new one for the VNN-LIB format, supporting the standardization effort of the community.
In addition to the aforementioned network and property file formats, Marabou also supports a native query file format that describes a set of linear and non-linear constraints. This can be dumped/parsed from all interfaces.
2.5 License and Installation
Marabou is available under the permissive modified BSD open-source license, and runs on Linux and macOS machines. The tool can be built from scratch using CMake. Marabou is now also available on The Python Package Index (PyPI) and can be installed through pip.
3 Highlighted Features and Applications
In terms of performance, Marabou is on par with state-of-the-art verification tools. In the latest VNN-COMP [11], Marabou won the second place overall, and scored the highest among all CPU-based verifiers. We summarize the main results in Appendix 0.B. In this section, we focus on the usability aspect of Marabou, and highlight some of its recent applications — as well as the features that make them possible. We believe this diverse set of use cases (as well as relevant scripts to be released in the artifact) serve as valuable examples, which will inspire new ways to apply the solver. A runtime evaluation of Marabou 2.0 against an early version appears in Section 4.
3.0.1 Verifying the Decima job scheduler.
Recently, Graph Neural Networks (GNNs) have been used to schedule jobs over multi-user, distributed-computing clusters, achieving state-of-the-art job completion time [47]. However, concerns remain over whether GNN-based solutions satisfy expected cost-critical properties beyond performance. Marabou has been used to verify a well-known fairness property called strategy-proofness [69] for a high-profile, state-of-the-art GNN-based scheduler called Decima [47]. The verified property states that “a user cannot get their job scheduled earlier by misrepresenting their resource requirement.” While it is challenging to represent a GNN directly in ONNX [25], Marabou’s Python API makes it possible to manually encode Decima and the specification as a set of linear and non-linear constraints. From these constraints, the Network-level Reasoner is able to automatically infer a feed-forward structure with residual connections and then use it for the purpose of abstract interpretation. Notably, Marabou was able to handle the original Decima architecture, proving that the property holds on the vast majority of the examined job profiles but can indeed be violated in some cases.
3.0.2 Formal XAI.
Despite their prevalence, DNNs are considered “black boxes”, uninterpretable to humans. Explainable AI (XAI) aims to understand DNN decisions to enhance trust. Most XAI methods are heuristic-based and lack formal correctness guarantees [53, 45, 54], which can be problematic for critical, regulation-heavy systems. Recent work showed that Marabou can be utilized as a sub-routine in procedures designed for producing formal and provable explanations for DNNs [71, 7, 8, 46, 32]. For instance, it can be used in constructing formal abductive explanations [35, 8], which are subsets of input features that are, by themselves, provably sufficient for determining the DNN’s output. This approach has been successfully applied to large DNNs in the domains of computer vision [71, 8], NLP [46], and DRL robotic navigation [7]. These studies highlight the potential of Marabou in tasks that go beyond formal verification.
3.0.3 Analyzing learning-based robotic systems.
Deep Reinforcement Learning has extensive application in robotic planning and control. Marabou has been applied in these settings to analyze different safety and liveness properties [17, 3, 2, 64]. For example, Amir et al. [2] used Marabou to detect infinite loops in a real-world robotic navigation platform. This was achieved by querying whether there exists a state to which the robot will always return within a finite number of steps $k$, effectively entering an infinite loop. A multi-step property like this can be conveniently encoded in Marabou, by (i) encoding $k$ copies of the control policy; (ii) for each time-step $t$, encoding the system transition as constraints over the current state (input to the policy at $t$), the decided action (output of the policy at $t$), and the next state (input to the policy at $t+1$); and (iii) encoding the “loop” constraint that the initial state ($t_{1}$) is equal to the final state ($t_{k}$). From this set of constraints, the Network-level Reasoner can infer the structure of and perform abstract interpretations over a concatenated network, where the input is the initial state and the output is the final state. Moreover, due to the low input dimension, the split-and-conquer mode in the Multi-thread Manager can be used to perform input-splitting, effectively searching for such loops in independent input regions in parallel. Notably, Marabou can detect loops in the system for agents trained using state-of-the-art RL algorithms, in cases where gradient/optimization-based approaches fail to find any. Loops detected this way have also been observed in the real world [1].
3.0.4 Proof production for the ACAS-Xu benchmarks.
A well-studied set of benchmarks in DNN verification derives from an implementation of the ACAS-Xu airborne system for collision avoidance [38]. Using Marabou, we were able to produce certificates of unsatisfiability for these benchmarks for the first time. Marabou was able to produce certificates for 113 out of the 180 tested benchmarks, with only mild overhead incurred by proof generation and certification. The proof certificates contained over 1.46 million proof-tree leaves, of which more than $99\%9$ were certified by the native proof checker, while the remaining were certified by a trusted SMT solver. Additional details are provided in [36].
3.0.5 Specifications on neural activation patterns.
Properties of hidden neurons garner increasing interest [75], as they shed light on the internal decision-making process of the neural network. Gopinath et al. [23] observed that for a fixed neural network, certain neuron activation patterns (NAPs) empirically entail a fixed prediction. More recently, Geng et al. [20] formally verified (using Marabou) the aforementioned property, along with a variety of other properties related to NAPs. Specifications related to NAPs can be conveniently encoded in Marabou. For example, specifying that a certain ReLU is activated amounts to setting the lower bound of the variable corresponding to the ReLU input to 0, using the general constraint-encoding methods in the Python/C++ API. Constraints on internal neurons, as with other constraints, can be propagated by the Preprocessor and Network-level Reasoner to tighten bounds.
3.0.6 Robustness against semantically meaningful perturbations.
Considering specifications of perception networks, there is an ongoing effort in the verification community to go beyond adversarial robustness [5, 41, 48, 70, 49]. Marabou has been used to verify robustness against semantically meaningful perturbations that can be analytically defined/abstracted as linear constraints on the neural network inputs (e.g., brightness, uniform haze) [51]. More recently, Marabou has also been successfully applied in a neural symbolic approach, where the correct network behavior is defined with respect to that of another network [72, 70]. For example, Wu et al. [70] considered the specification that an image classifier’s prediction does not change with respect to outputs of an image generative model trained to capture a complex distribution shift (e.g., change in weather condition). A property like this can be conveniently defined in Marabou by loading the classifier and the generator through the Python API and adding the relevant constraints on/between their input and output variables.
3.0.7 Minimal modifications of neural networks.
Recently, multiple papers have proposed techniques to repair a neural network (e.g., by altering selected weights or adding patches) against input points that result in buggy behaviors. Marabou has also been applied in this context to obtain provably minimal modification to the DNN’s weights so that a previously mispredicted input would become correctly predicted. We refer the reader to [52, 22] for additional details.
3.0.8 Verification of quantized neural Networks.
Quantization [37] has become a standard technique for deploying deep neural networks on resource-constrained hardware. The process involves replacing floating point arithmetic with integer arithmetic. Recently, Huang, et.al. [31] have leveraged Marabou to verify quantized neural networks (QNNs). While mainstream platforms like PyTorch do not support saving QNN models in the standard ONNX format, Marabou’s flexible Python API makes it possible to manually encode the quantization operations with constraints such as Round and Clip.
3.0.9 Pruning and slicing neural networks.
Another well-studied approach to reduce the computational cost of deep neural networks is pruning [28, 34], which involves identifying and removing neurons that contribute minimally to the network’s output. Most existing approaches are heuristic-based and lack formal guarantees that the simplified network preserves the behavior of the original. In contrast, it has been shown that Marabou can be used as a sub-procedure in a pruning algorithm [21, 42], which is able to reduce the size of DNNs while provably preserving the output of the neural network up to a given error margin.
3.0.10 Integrating Marabou into complex system verification.
There is a growing demand to embed neural network verification into verification of complex systems that rely on neural networks for perception or decision-making [50]. There are multiple challenges in verifying such systems, and we mention two of them here: (i) currently available provers struggle to overcome the embedding gap, which arises because complex system properties are stated using the abstractions of the problem domain, whereas neural network propeties are stated at a low-level, in terms of normalized real vectors; and (ii) more generally, higher-order theorem provers like HOL, Coq, and Agda can prove more abstract properties of complex systems, but fail at neural network verification; and vice versa: neural network verifiers are not designed to reason about complex hybrid systems. The domain-specific language Vehicle [15, 14] aims at bridging this gap; and includes Marabou as its default verifier. This choice was made thanks to Marabou’s syntax, which is richer and more flexible than available in other, more specialized neural network verifiers.
4 Runtime Evaluation
Figure 2: Runtime performance of Marabou 2.0 and an early version of Marabou on four applications supported by both versions.
We measure the performance improvement in Marabou 2.0 by comparing it against an early Marabou version (git commit 1c1c66), which can handle ReLU and Max constraints and supports symbolic bound propagation [65]. We collected four benchmark sets from the applications described in Section  3: Alternating Loop [2], DeepCert [51], NAP [23, 20], and VeriX [71]. There are 745 instances in total. Details about the benchmarks can be found in Appendix 0.C.1.
Figure 2 compares the runtime of the two Marabou versions on all the benchmarks with a 1 hour CPU timeout. Each configuration was given 1 core and 8GB of memory. Note that Marabou 2.0 was not configured with external solvers in this experiment. We see that Marabou 2.0 is significantly more efficient for a vast majority of the instances. Upon closer examination, an at-least 2×2$\times$2 × speed-up is achieved on 417 instances and an at-least 10×10$\times$10 × speed-up is achieved on 253 instances. Moreover, Marabou 2.0 is also significantly more memory efficient, with a median peak usage of 59MB (versus 604MB with the old version). Solvers’performance on individual benchmarks is reported in Appendix 0.C.2.
5 Conclusion and Next Steps
We have summarized the current state of Marabou, a maturing formal analyzer for neural-network-enabled systems that is under active development. In its current form, Marabou is a versatile and user-friendly toolkit suitable for a wide range of formal analysis tasks. Moving forward, we plan to improve Marabou in several dimensions. Currently, we are actively integrating a CDCL mechanism in the SMT Solver module. Given that many applications involve repeated invocation of the solver on similar queries, we also plan to support incremental solving in the style of pushing and popping constraints, leveraging the newly introduced context-dependent data structures. In addition, adding GPU support (in the Network-level Reasoner) and handling other types of non-linear constraints are also on the development agenda for Marabou.
5.0.1 Acknowledgment
The work of Wu, Zeljić, Tagomori, Huang and Wu was partially supported by the NSF (grant number 2211505), by the BSF (grant number 2020250), a Ford Alliance Project (199909), the Stanford Center for AI Safety, and the Stanford Institute for Human-Centered Artificial Intelligence (HAI). The work of Daggit, Kokke and Komendantskaya was partially supported by the EPSRC grant EP/T026952/1, AISEC: AI Secure and Explainable by Construction. The work of Isac, Refaeli, Amir, Bassan, Lahav and Katz was partially funded by the ISF (grant number 3420/21), by the BSF (grant numbers 2021769 and 2020250), and by the European Union (ERC, VeriDeL, 101112713). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of Zhang was partially supported by the NSFC (grant number 62161146001).
Appendix 0.A Architecture and Core Components
0.A.1 Supported Non-linear Constraints in Marabou 2.0
Type Preproc. NLR SMT (MI)LP CEGAR Proof
ReLU
Max
DNF
Sign
Absolute value
Leaky ReLU
Round
Clip
Sigmoid
Tanh
Softmax
Bi-linear
0.A.2 Pythonic Interface Example
We provide an example to illustrate the newly introduced Pythonic interface. In the example, we check whether the first output of a network (stored in the ONNX format) can be less than or equal to half of its second output, when the first input is greater than or equal to 0.10.10.1. The base Python API to implement this is shown in (a), and (b) exhibits the “Pythonic” API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.setLowerBound(X[0], 0.1)
Q.addInequality([Y[0], Y[1]], [1, -0.5], 0)
Q.solve()
((a)) The base Python API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.addConstraint(Var(X[0]) >= 0.1)
Q.addConstraint(Var(Y[0]) <= 0.5 * Var(Y[1]))
Q.solve()
((b)) The “Pythonic” API.
Appendix 0.B Marabou in VNN-COMP 2023 [11]
Marabou participated in the Fourth International Verification of Neural Networks Competition (VNN-COMP’23) and won the second place overall and scored the highest among all CPU-based verifiers. Table 1 shows the total number of instances solved by the participating solvers. Detailed information can be found in the VNN-COMP 2023 report [11]. The participating version is publicly available in the Marabou repo (commit 1a3ca6). Overall, on the competition benchmarks, there is a non-trivial performance gap between Marabou and GPU-based solvers like $\alpha$-$\beta$-CROWN, especially for larger networks. Supporting GPUs in the Network-level Reasoner is on the development agenda for Marabou.
Table 1: Total number of solved instances in VNN-COMP’23
# Tool Count GPU Support
1 $\alpha$-$\beta$-CROWN 721
2 Marabou 594
3 NeuralSAT 452
4 PyRAT 416
5 nnenum 319
6 NNV 155
7 FastBATLLNN 32
Appendix 0.C Runtime Evaluation
0.C.1 Descriptions of the Benchmarks
The DeepCert benchmarks come from a study [51] of the robustness of image classifiers against a set of analytically defined contextually relevant image perturbations. In particular, we focus on the haze perturbation on an image classifier (model 2A) for the GTSRB dataset. The Alternating Loop benchmarks come from a study of the existence of infinite loops in a robot navigation system [2]. The NAP benchmarks are less straightforward to procure: each benchmark considers an NAP mined on one of the ACAS-Xu networks by Gopinath et al. [23] and checks the robustness of that NAP as defined by Geng et al. [20]. We hope to also evaluate Marabou on the NAP robustness benchmarks described in the latter work, but they are not yet publicly available. Finally, the VeriX benchmarks involve performing a formal explanation extraction procedure (which repeatedly queries Marabou) on a MNIST classifier (the same one used in the original work).
0.C.2 Solver Performance on Individual Benchmarks
Marabou 2.0 Marabou (1c1c66)
Family Benchmark Result Time Mem. Result Time Mem.
[…]
Alt. Loops REI.id114.ep84297 UNS 54.1 9.6 UNS 1036.7 50.9
Alt. Loops REI.id114.ep94159 SAT 0.6 6.9 SAT 22.3 42.6
Alt. Loops REI.id114.ep94581 SAT 0.8 6.9 SAT 35.9 46.2
Alt. Loops REI.id114.ep97993 SAT 0.7 7.1 SAT 139.3 48.4
Alt. Loops REI.id128.ep68334 UNS 116.6 10.1 UNS 1749.2 56.0
Alt. Loops REI.id128.ep85878 UNS 168.6 10.6 UNS 3478.2 60.4
Alt. Loops REI.id128.ep87384 SAT 0.6 6.6 SAT 308.5 57.1
Alt. Loops REI.id128.ep91042 UNS 85.5 10.3 UNS 1267.4 53.9
Alt. Loops REI.id128.ep97983 UNS 147.3 10.7 UNS 2404.6 56.7
Alt. Loops REI.id158.ep53113 SAT 25.9 10.6 SAT 16.9 34.8
Alt. Loops REI.id158.ep53499 SAT 19.4 10.6 SAT 34.1 36.8
Alt. Loops REI.id158.ep53554 SAT 12.9 10.5 SAT 17.7 39.0
Alt. Loops REI.id158.ep59378 UNS 33.3 9.3 UNS 80.8 37.6
Alt. Loops REI.id158.ep63427 UNS 28.3 9.9 UNS 68.9 38.7
Alt. Loops REI.id165.ep82822 SAT 0.7 7.0 SAT 13.2 31.6
Alt. Loops REI.id165.ep94134 SAT 0.7 7.0 SAT 14.4 33.9
Alt. Loops REI.id165.ep94522 SAT 1.0 7.1 SAT 0.2 11.2
Alt. Loops REI.id165.ep95976 SAT 0.7 7.0 SAT 30.7 34.6
Alt. Loops REI.id165.ep96974 SAT 35.8 9.7 SAT 16.8 32.5
Alt. Loops REI.id174.ep66258 SAT 16.8 10.6 SAT 0.1 5.8
Alt. Loops REI.id174.ep81057 SAT 0.6 7.0 SAT 0.3 19.6
Alt. Loops REI.id174.ep91107 SAT 25.6 10.5 SAT 49.7 43.4
Alt. Loops REI.id174.ep97933 SAT 0.7 6.8 SAT 10.6 41.8
Alt. Loops REI.id174.ep98455 SAT 21.5 10.9 SAT 0.2 13.0
Alt. Loops REI.id176.ep92206 UNS 635.3 11.0 UNK 3601.0 66.0
Alt. Loops REI.id176.ep95516 UNS 375.7 10.6 UNK 3601.0 61.5
Alt. Loops REI.id176.ep96138 UNS 340.2 10.6 UNK 3601.0 63.1
Alt. Loops REI.id176.ep98430 UNS 548.5 10.9 UNK 3601.0 64.6
Alt. Loops REI.id176.ep99427 UNS 536.1 10.7 UNK 3601.0 64.2
Alt. Loops REI.id180.ep58329 SAT 0.5 6.7 SAT 0.3 15.8
Alt. Loops REI.id180.ep58461 SAT 0.7 6.7 SAT 0.2 14.3
Alt. Loops REI.id180.ep61293 SAT 0.8 7.0 SAT 0.6 28.1
Alt. Loops REI.id180.ep69795 SAT 0.7 6.9 SAT 0.2 9.6
Alt. Loops REI.id180.ep77624 SAT 0.7 7.2 SAT 0.1 9.3
Alt. Loops REI.id196.ep41633 SAT 0.9 7.0 SAT 0.2 12.0
Alt. Loops REI.id196.ep43251 SAT 0.8 7.0 SAT 0.1 9.3
Alt. Loops REI.id196.ep73462 SAT 0.7 6.9 SAT 0.2 12.8
Alt. Loops REI.id196.ep86142 SAT 0.8 7.1 SAT 0.1 8.3
Alt. Loops REI.id196.ep99092 SAT 0.6 6.9 SAT 24.0 34.0
Alt. Loops REI.id201.ep39282 SAT 17.4 10.1 SAT 4.0 39.8
Alt. Loops REI.id201.ep48186 SAT 0.7 7.0 SAT 51.4 43.1
Alt. Loops REI.id201.ep48219 SAT 0.6 6.6 SAT 0.2 10.6
Alt. Loops REI.id201.ep56910 SAT 0.8 7.1 SAT 28.7 34.9
Alt. Loops REI.id201.ep88825 SAT 0.6 6.5 SAT 49.6 45.4
Alt. Loops REI.id215.ep96277 SAT 0.3 6.6 SAT 1.7 41.1
Alt. Loops REI.id215.ep97333 SAT 0.4 6.3 SAT 4.4 40.8
Alt. Loops REI.id215.ep97355 SAT 0.3 6.6 SAT 0.6 32.9
Alt. Loops REI.id215.ep97421 SAT 0.7 6.6 SAT 1763.0 60.1
Alt. Loops REI.id215.ep98767 SAT 0.6 6.9 SAT 2.1 35.0
Alt. Loops REI.id234.ep70556 SAT 0.7 6.8 SAT 0.1 10.8
Alt. Loops REI.id234.ep71035 SAT 0.6 6.6 SAT 0.3 24.4
Alt. Loops REI.id234.ep71965 SAT 0.7 7.0 SAT 0.2 11.8
Alt. Loops REI.id234.ep83703 SAT 0.5 6.9 SAT 0.3 22.4
Alt. Loops REI.id234.ep83780 SAT 0.7 8.0 SAT 1.0 33.8
Alt. Loops REI.id239.ep71386 UNS 197.2 10.3 UNS 1691.7 55.5
Alt. Loops REI.id239.ep73100 UNS 186.9 10.6 UNS 1110.3 60.0
Alt. Loops REI.id239.ep73504 SAT 59.0 11.1 SAT 3.4 51.2
Alt. Loops REI.id239.ep79112 UNS 190.0 10.4 UNS 1643.7 56.7
Alt. Loops REI.id239.ep90118 SAT 30.8 9.9 SAT 0.4 18.0
Alt. Loops REI.id240.ep85050 UNS 40.3 10.0 UNS 308.5 47.0
Alt. Loops REI.id240.ep89317 UNS 32.5 10.7 UNS 152.7 42.7
Alt. Loops REI.id240.ep91714 UNS 26.8 9.8 UNS 177.1 42.5
Alt. Loops REI.id240.ep98115 UNS 40.2 10.7 UNS 282.4 47.6
Alt. Loops REI.id240.ep98891 UNS 31.9 10.2 UNS 76.3 45.4
Alt. Loops REI.id267.ep91426 UNS 0.9 6.5 UNS 0.7 14.5
Alt. Loops REI.id267.ep95713 UNS 1.1 6.5 UNS 2.5 16.7
Alt. Loops REI.id267.ep96300 UNS 0.9 6.8 UNS 1.0 14.8
Alt. Loops REI.id267.ep96333 UNS 1.4 7.0 UNS 0.5 12.4
Alt. Loops REI.id267.ep96426 UNS 1.5 6.9 UNS 1.5 15.5
Alt. Loops REI.id286.ep89988 UNS 38.1 9.1 UNS 41.8 35.9
Alt. Loops REI.id286.ep89999 UNS 38.4 9.1 UNS 41.7 35.8
Alt. Loops REI.id286.ep90032 UNS 32.5 9.7 UNS 59.7 37.9
Alt. Loops REI.id286.ep94925 UNS 34.0 9.0 UNS 56.8 38.1
Alt. Loops REI.id286.ep95719 UNS 27.6 9.4 UNS 63.4 38.2
Alt. Loops REI.id296.ep74854 UNS 127.3 11.1 UNS 697.8 51.9
Alt. Loops REI.id296.ep74969 UNS 108.8 10.0 UNS 559.8 55.9
Alt. Loops REI.id296.ep76414 UNS 113.8 10.3 UNS 500.1 51.7
Alt. Loops REI.id296.ep77867 UNS 116.8 10.0 UNS 131.1 48.4
Alt. Loops REI.id296.ep79125 UNS 111.0 10.4 UNS 274.0 46.4
Alt. Loops REI.id298.ep75768 UNS 167.3 9.8 UNK 3601.0 66.3
Alt. Loops REI.id298.ep77187 SAT 0.8 6.9 SAT 0.4 36.5
Alt. Loops REI.id298.ep79143 UNS 162.0 10.3 UNK 3601.0 70.0
Alt. Loops REI.id298.ep85589 UNS 198.8 11.0 UNK 3601.0 65.9
Alt. Loops REI.id298.ep97154 SAT 0.6 7.0 SAT 0.1 8.0
Alt. Loops REI.id299.ep69103 UNS 122.0 11.0 UNS 479.8 49.8
Alt. Loops REI.id299.ep69400 UNS 98.1 10.9 UNS 361.7 47.0
Alt. Loops REI.id299.ep81175 UNS 79.2 10.6 UNS 284.3 45.7
Alt. Loops REI.id299.ep92705 UNS 90.0 10.6 UNS 739.8 50.7
Alt. Loops REI.id299.ep99426 UNS 49.8 10.3 UNS 569.1 49.5
Alt. Loops REI.id303.ep68520 UNS 402.2 12.0 UNK 3601.0 64.6
Alt. Loops REI.id303.ep69708 SAT 1.7 7.6 SAT 987.6 61.3
Alt. Loops REI.id303.ep94994 SAT 0.7 6.6 SAT 72.8 58.6
Alt. Loops REI.id303.ep95038 SAT 0.9 7.1 SAT 4.4 52.3
Alt. Loops REI.id303.ep95137 SAT 1.4 7.2 SAT 1399.1 72.3
Alt. Loops REI.id308.ep94787 UNS 39.2 9.8 UNS 95.0 46.1
Alt. Loops REI.id308.ep97397 SAT 0.7 7.3 SAT 0.2 14.4
Alt. Loops REI.id308.ep98363 SAT 0.6 7.3 SAT 125.7 45.1
Alt. Loops REI.id308.ep98830 UNS 47.3 9.8 UNS 207.0 46.9
Alt. Loops REI.id308.ep99303 SAT 1.0 6.7 SAT 6.1 35.9
Alt. Loops REI.id318.ep75036 SAT 0.4 6.6 SAT 14.4 64.8
Alt. Loops REI.id318.ep85341 SAT 50.0 10.4 SAT 164.6 57.5
Alt. Loops REI.id318.ep89536 SAT 0.8 7.0 SAT 27.3 55.8
Alt. Loops REI.id318.ep94480 SAT 0.7 6.9 SAT 0.9 37.7
Alt. Loops REI.id318.ep96308 SAT 0.7 7.0 SAT 0.4 21.7
Alt. Loops REI.id319.ep62661 SAT 0.8 6.9 SAT 71.0 54.8
Alt. Loops REI.id319.ep84434 SAT 0.6 7.0 SAT 2437.7 67.5
Alt. Loops REI.id319.ep84478 SAT 0.8 7.0 SAT 0.3 14.3
Alt. Loops REI.id319.ep86630 SAT 0.8 7.0 SAT 573.1 62.2
Alt. Loops REI.id319.ep89649 SAT 0.7 6.7 SAT 0.3 12.9
Alt. Loops REI.id321.ep55647 UNS 9.1 9.2 UNS 19.9 33.0
Alt. Loops REI.id321.ep81083 UNS 34.1 8.8 UNS 404.1 53.6
Alt. Loops REI.id321.ep86834 UNS 25.6 9.2 UNS 712.3 62.0
Alt. Loops REI.id321.ep92994 UNS 48.9 9.9 UNS 389.5 54.3
Alt. Loops REI.id321.ep93439 UNS 24.8 9.9 UNS 686.8 58.2
Alt. Loops REI.id343.ep55239 UNS 83.2 9.2 UNS 149.7 47.1
Alt. Loops REI.id343.ep63040 UNS 77.5 9.8 UNS 168.2 45.4
Alt. Loops REI.id343.ep63051 UNS 76.7 10.3 UNS 166.6 45.5
Alt. Loops REI.id343.ep73171 UNS 160.2 10.2 UNS 1022.7 54.2
Alt. Loops REI.id343.ep96195 UNS 241.5 9.9 UNS 3113.7 57.7
Alt. Loops REI.id35.ep61486 SAT 0.4 6.3 SAT 22.1 39.1
Alt. Loops REI.id35.ep61788 SAT 0.4 6.5 SAT 0.2 10.9
Alt. Loops REI.id35.ep81067 SAT 0.6 7.2 SAT 0.2 12.3
Alt. Loops REI.id35.ep85842 SAT 0.6 6.6 SAT 0.2 10.4
Alt. Loops REI.id35.ep94624 SAT 0.3 6.4 SAT 0.1 7.1
Alt. Loops REI.id352.ep49376 UNS 8.6 9.1 UNS 58.8 39.3
Alt. Loops REI.id352.ep51510 UNS 12.2 8.6 UNS 169.9 43.4
Alt. Loops REI.id352.ep68171 UNS 13.4 9.6 UNS 87.1 39.8
Alt. Loops REI.id352.ep69872 UNS 10.8 9.4 UNS 68.1 38.3
Alt. Loops REI.id352.ep70312 UNS 12.6 9.2 UNS 24.0 37.8
Alt. Loops REI.id379.ep71969 SAT 3.8 9.8 SAT 13.7 43.2
Alt. Loops REI.id379.ep80359 SAT 0.6 6.7 SAT 0.2 11.7
Alt. Loops REI.id379.ep87708 UNS 214.8 11.1 UNS 2425.8 57.3
Alt. Loops REI.id379.ep96909 SAT 1.9 8.7 SAT 180.9 69.4
Alt. Loops REI.id379.ep97133 SAT 0.5 6.7 SAT 0.7 36.4
Alt. Loops REI.id381.ep85470 SAT 58.1 12.1 UNK 3601.0 64.0
Alt. Loops REI.id381.ep85558 SAT 0.7 6.7 SAT 1140.9 62.2
Alt. Loops REI.id381.ep92541 UNS 321.6 10.9 UNK 3600.9 59.8
Alt. Loops REI.id381.ep92994 SAT 1.6 7.4 UNK 3601.0 64.2
Alt. Loops REI.id381.ep93562 UNS 315.7 11.5 UNK 3601.0 64.3
Alt. Loops REI.id393.ep88789 SAT 18.4 11.9 SAT 0.3 17.0
Alt. Loops REI.id393.ep90550 SAT 0.8 7.2 SAT 64.9 56.5
Alt. Loops REI.id393.ep97111 SAT 19.3 11.6 SAT 0.1 10.0
Alt. Loops REI.id393.ep97199 SAT 23.3 11.3 SAT 0.2 15.2
Alt. Loops REI.id393.ep97221 SAT 23.8 10.4 SAT 0.2 14.3
Alt. Loops REI.id399.ep62611 UNS 45.5 9.7 UNS 490.9 56.3
Alt. Loops REI.id399.ep69315 UNS 56.0 10.4 UNS 3342.7 68.6
Alt. Loops REI.id399.ep85537 UNS 76.4 9.5 UNS 2220.3 60.8
Alt. Loops REI.id399.ep92145 UNS 69.7 10.6 UNS 1475.9 64.6
Alt. Loops REI.id399.ep96449 SAT 1.1 7.1 SAT 140.4 64.8
Alt. Loops REI.id418.ep91524 SAT 0.8 6.9 SAT 5.3 53.1
Alt. Loops REI.id418.ep92561 SAT 1.2 7.1 SAT 19.8 51.7
Alt. Loops REI.id418.ep92671 SAT 1.5 7.4 SAT 30.0 49.7
Alt. Loops REI.id418.ep99126 SAT 96.6 11.0 SAT 134.2 51.4
Alt. Loops REI.id418.ep99947 SAT 0.7 7.1 SAT 13.4 51.0
Alt. Loops REI.id425.ep69300 SAT 0.7 6.6 SAT 0.8 37.0
Alt. Loops REI.id425.ep83340 SAT 0.3 6.5 SAT 0.3 14.0
Alt. Loops REI.id425.ep93745 UNS 166.7 11.0 UNS 1299.1 51.4
Alt. Loops REI.id425.ep96981 SAT 0.7 6.9 SAT 0.1 8.0
Alt. Loops REI.id425.ep98177 SAT 0.8 7.1 SAT 0.1 11.8
Alt. Loops REI.id444.ep77347 UNS 90.0 10.3 UNS 501.8 50.7
Alt. Loops REI.id444.ep78487 UNS 89.7 11.1 UNS 1094.5 51.6
Alt. Loops REI.id444.ep87148 UNS 71.6 9.8 UNS 842.5 51.1
Alt. Loops REI.id444.ep92635 UNS 96.3 10.8 UNK 3601.0 48.8
Alt. Loops REI.id444.ep96160 UNS 95.2 10.1 UNS 1262.5 51.4
Alt. Loops REI.id457.ep46821 SAT 9.4 8.8 SAT 0.1 8.8
Alt. Loops REI.id457.ep69348 SAT 0.4 6.5 SAT 0.2 10.6
Alt. Loops REI.id457.ep78403 SAT 0.7 7.1 SAT 0.1 7.3
Alt. Loops REI.id457.ep79411 SAT 0.6 6.7 SAT 0.2 12.7
Alt. Loops REI.id457.ep81404 SAT 0.5 6.8 SAT 0.1 8.5
Alt. Loops REI.id47.ep79665 SAT 1.1 7.0 SAT 0.3 19.2
Alt. Loops REI.id47.ep92041 SAT 0.6 6.7 SAT 0.1 10.0
Alt. Loops REI.id47.ep96122 SAT 0.3 6.5 SAT 0.4 21.0
Alt. Loops REI.id47.ep99600 SAT 0.2 6.5 SAT 0.3 18.8
Alt. Loops REI.id47.ep99688 SAT 0.5 6.6 SAT 0.3 18.2
Alt. Loops REI.id491.ep54571 SAT 0.6 7.0 SAT 0.6 25.5
Alt. Loops REI.id491.ep67504 SAT 0.6 6.7 SAT 1.3 33.2
Alt. Loops REI.id491.ep67924 SAT 0.6 6.6 SAT 0.2 12.8
Alt. Loops REI.id491.ep70134 SAT 0.6 7.0 SAT 0.4 23.2
Alt. Loops REI.id491.ep96326 SAT 0.7 6.8 SAT 5.6 39.0
Alt. Loops REI.id502.ep71027 SAT 0.6 6.9 SAT 56.7 42.9
Alt. Loops REI.id502.ep71445 SAT 0.6 6.9 SAT 15.9 46.3
Alt. Loops REI.id502.ep84337 SAT 0.5 6.7 SAT 2.1 30.2
Alt. Loops REI.id502.ep86449 SAT 0.9 6.8 SAT 21.3 42.0
Alt. Loops REI.id502.ep90122 SAT 0.8 6.9 SAT 266.6 45.8
Alt. Loops REI.id512.ep70217 SAT 0.4 6.9 SAT 1.3 41.2
Alt. Loops REI.id512.ep70283 SAT 0.7 6.7 SAT 1.0 33.3
Alt. Loops REI.id512.ep75804 SAT 0.7 7.1 SAT 986.8 50.4
Alt. Loops REI.id512.ep97330 SAT 0.7 7.0 SAT 0.1 11.2
Alt. Loops REI.id512.ep97385 SAT 0.7 6.7 SAT 567.2 47.5
Alt. Loops REI.id518.ep68431 UNS 101.2 9.2 UNS 523.8 46.8
Alt. Loops REI.id518.ep70223 UNS 80.0 9.4 UNS 1358.8 49.7
Alt. Loops REI.id518.ep95639 UNS 126.9 10.7 UNS 132.8 39.7
Alt. Loops REI.id518.ep95900 UNS 112.7 10.2 UNS 602.4 45.1
Alt. Loops REI.id518.ep96115 SAT 83.8 10.2 UNS 561.7 45.2
Alt. Loops REI.id528.ep55711 SAT 1.4 7.4 SAT 1446.3 51.4
Alt. Loops REI.id528.ep89177 SAT 0.8 7.0 SAT 53.8 53.1
Alt. Loops REI.id528.ep89331 SAT 0.8 6.7 SAT 158.9 48.1
Alt. Loops REI.id528.ep90341 SAT 0.7 6.5 SAT 36.3 39.8
Alt. Loops REI.id528.ep95627 SAT 0.7 6.9 SAT 0.1 8.3
Alt. Loops REI.id530.ep82130 SAT 0.4 6.5 SAT 0.2 13.2
Alt. Loops REI.id530.ep82185 SAT 0.7 7.4 SAT 0.3 12.9
Alt. Loops REI.id530.ep82207 SAT 0.5 6.7 SAT 0.2 13.1
Alt. Loops REI.id530.ep91636 SAT 0.6 7.0 SAT 0.4 31.3
Alt. Loops REI.id530.ep94463 SAT 0.5 6.9 SAT 0.2 12.8
Alt. Loops REI.id535.ep76894 UNS 322.8 9.7 UNS 54.5 43.6
Alt. Loops REI.id535.ep76916 UNS 23.1 9.8 UNS 51.4 39.6
Alt. Loops REI.id535.ep87193 SAT 0.8 7.0 SAT 0.2 13.9
Alt. Loops REI.id535.ep87226 SAT 0.9 7.1 SAT 0.5 22.8
Alt. Loops REI.id535.ep91323 SAT 25.2 9.4 UNS 176.3 44.0
Alt. Loops REI.id540.ep57595 UNS 1472.6 10.9 UNK 3601.0 65.2
Alt. Loops REI.id540.ep57610 UNS 1646.9 11.0 UNK 3601.0 66.9
Alt. Loops REI.id540.ep57775 UNS 2759.4 11.2 UNK 3601.0 68.2
Alt. Loops REI.id540.ep85904 SAT 0.9 7.0 SAT 762.1 68.3
Alt. Loops REI.id540.ep94690 SAT 0.8 7.1 SAT 0.1 12.2
Alt. Loops REI.id549.ep85789 SAT 0.7 6.9 SAT 0.2 15.2
Alt. Loops REI.id549.ep87887 SAT 510.8 11.4 SAT 1509.6 64.1
Alt. Loops REI.id549.ep93366 SAT 0.9 6.9 SAT 28.6 51.1
Alt. Loops REI.id549.ep99098 SAT 0.9 6.9 SAT 0.3 16.4
Alt. Loops REI.id549.ep99331 SAT 0.7 6.7 SAT 0.1 8.2
Alt. Loops REI.id67.ep82981 UNS 26.2 11.2 UNS 48.7 33.4
Alt. Loops REI.id67.ep84585 UNS 54.5 10.7 UNS 98.8 39.6
Alt. Loops REI.id67.ep87611 UNS 53.7 10.9 UNS 29.7 34.7
Alt. Loops REI.id67.ep95115 SAT 0.6 6.8 SAT 111.2 41.5
Alt. Loops REI.id67.ep99214 SAT 1.0 7.1 SAT 0.8 35.5
Alt. Loops REI.id68.ep95777 SAT 0.6 7.0 SAT 0.4 20.0
Alt. Loops REI.id68.ep95821 SAT 0.5 6.9 SAT 0.2 10.8
Alt. Loops REI.id68.ep95854 SAT 0.5 6.6 SAT 0.3 16.7
Alt. Loops REI.id68.ep95942 SAT 0.4 6.7 SAT 0.2 11.4
Alt. Loops REI.id68.ep98804 SAT 0.8 7.3 SAT 0.3 23.0
Alt. Loops REI.id73.ep74414 SAT 299.2 11.7 SAT 81.0 52.7
Alt. Loops REI.id73.ep74469 SAT 262.1 11.9 SAT 23.7 50.4
Alt. Loops REI.id73.ep82370 UNS 325.1 11.7 UNS 1166.3 58.1
Alt. Loops REI.id73.ep84023 UNS 296.5 11.2 UNS 946.3 54.9
Alt. Loops REI.id73.ep91504 SAT 0.6 6.8 SAT 153.7 53.7
Alt. Loops REI.id76.ep90543 UNS 288.1 10.6 UNS 2095.2 62.3
Alt. Loops REI.id76.ep91480 UNS 255.0 11.3 UNS 1264.9 62.4
Alt. Loops REI.id76.ep92143 UNS 289.1 10.5 UNS 411.0 46.4
Alt. Loops REI.id76.ep98692 UNS 270.3 12.0 UNS 1011.0 54.6
Alt. Loops REI.id76.ep99198 UNS 342.9 11.7 UNS 1862.4 56.5
Alt. Loops REI.id77.ep49509 UNS 39.8 9.3 UNS 69.4 37.8
Alt. Loops REI.id77.ep75183 UNS 38.7 9.2 UNS 102.2 38.4
Alt. Loops REI.id77.ep91446 UNS 49.6 9.6 UNS 165.0 41.0
Alt. Loops REI.id77.ep99119 UNS 67.1 9.4 UNS 147.6 38.8
Alt. Loops REI.id77.ep99229 UNS 68.9 10.0 UNS 144.3 40.5
Alt. Loops REI.id81.ep89983 UNS 209.4 11.2 UNK 3601.0 59.4
Alt. Loops REI.id81.ep91408 UNS 204.4 10.7 UNK 3601.0 60.9
Alt. Loops REI.id81.ep91571 UNS 220.4 10.8 UNS 2485.5 56.8
Alt. Loops REI.id81.ep97775 UNS 313.1 11.4 UNS 3578.0 61.4
Alt. Loops REI.id81.ep97833 UNS 199.5 11.1 UNK 3601.0 57.0
Alt. Loops REI.id90.ep84292 UNS 2081.5 9.9 UNK 3601.0 60.4
Alt. Loops REI.id90.ep94747 UNS 246.2 11.2 UNK 3601.0 59.0
Alt. Loops REI.id90.ep97680 UNS 194.2 10.2 UNK 3601.0 54.8
Alt. Loops REI.id90.ep97955 UNS 219.2 10.9 UNK 3601.0 59.9
Alt. Loops REI.id90.ep99033 UNS 244.9 11.3 UNK 3601.0 58.0
Alt. Loops REI.id91.ep54663 UNS 43.6 9.6 UNS 1790.1 53.0
Alt. Loops REI.id91.ep58150 UNS 37.9 9.0 UNK 3601.0 55.4
Alt. Loops REI.id91.ep58172 UNS 32.6 8.7 UNS 2933.8 56.0
Alt. Loops REI.id91.ep75275 UNS 133.9 10.1 UNS 649.1 48.4
Alt. Loops REI.id91.ep95105 UNS 172.4 10.1 UNS 1257.0 56.5
DeepCert ind0.target6.eps0.2 UNS 10.4 412.4 UNS 268.9 603.7
DeepCert ind0.target6.eps0.4 UNS 10.3 425.5 UNS 830.0 2053.0
DeepCert ind0.target6.eps0.6 UNS 25.1 582.4 UNS 770.8 1252.2
DeepCert ind14.target6.eps0.2 UNS 10.3 412.9 UNS 457.9 1081.1
DeepCert ind14.target6.eps0.4 SAT 26.9 572.0 SAT 617.3 1421.2
DeepCert ind14.target6.eps0.6 SAT 25.2 580.8 SAT 428.7 1086.1
DeepCert ind15.target6.eps0.2 UNS 10.2 410.0 UNS 398.0 602.4
DeepCert ind15.target6.eps0.4 UNS 38.0 571.3 UNS 476.5 951.0
DeepCert ind15.target6.eps0.6 SAT 24.5 579.3 SAT 177.8 753.4
DeepCert ind19.target6.eps0.2 UNS 10.7 409.2 UNS 442.5 938.9
DeepCert ind19.target6.eps0.4 UNS 30.6 573.9 UNS 532.1 1419.9
DeepCert ind19.target6.eps0.6 SAT 26.4 582.6 SAT 419.0 754.9
DeepCert ind20.target6.eps0.2 UNS 10.6 413.6 UNS 352.1 603.2
DeepCert ind20.target6.eps0.4 UNS 26.6 567.3 UNS 1568.0 2244.6
DeepCert ind20.target6.eps0.6 SAT 24.6 579.8 SAT 436.5 1403.0
DeepCert ind21.target6.eps0.2 UNS 11.0 412.6 UNS 366.8 601.6
DeepCert ind21.target6.eps0.4 UNS 30.2 571.0 UNS 1054.4 1889.6
DeepCert ind21.target6.eps0.6 SAT 27.0 582.3 SAT 874.6 2369.0
DeepCert ind22.target6.eps0.2 UNS 10.6 408.4 UNS 261.3 602.2
DeepCert ind22.target6.eps0.4 UNS 26.0 567.8 UNS 698.6 943.7
DeepCert ind22.target6.eps0.6 SAT 26.5 576.0 SAT 376.1 3149.2
DeepCert ind23.target6.eps0.2 UNS 10.8 412.4 UNS 615.3 1075.1
DeepCert ind23.target6.eps0.4 UNS 32.7 573.7 UNS 727.4 1255.8
DeepCert ind23.target6.eps0.6 SAT 24.9 579.4 SAT 162.7 910.2
DeepCert ind24.target6.eps0.2 UNS 10.5 411.7 UNS 829.5 1259.0
DeepCert ind24.target6.eps0.4 UNS 40.3 569.8 UNS 1477.3 1929.0
DeepCert ind24.target6.eps0.6 SAT 26.2 574.3 SAT 272.1 933.5
DeepCert ind25.target6.eps0.2 UNS 11.0 413.6 UNS 933.3 1572.4
DeepCert ind25.target6.eps0.4 SAT 25.8 571.3 SAT 293.6 925.2
DeepCert ind25.target6.eps0.6 SAT 24.9 580.8 SAT 357.4 1251.2
DeepCert ind26.target6.eps0.2 UNS 10.6 413.0 UNS 515.5 781.7
DeepCert ind26.target6.eps0.4 UNS 25.1 571.1 UNS 1136.0 1887.4
DeepCert ind26.target6.eps0.6 SAT 25.1 579.7 SAT 1453.1 1568.0
DeepCert ind27.target6.eps0.2 UNS 10.3 408.0 UNS 500.2 781.4
DeepCert ind27.target6.eps0.4 SAT 24.5 571.3 SAT 628.1 1271.7
DeepCert ind27.target6.eps0.6 SAT 25.2 579.6 SAT 905.0 2196.7
DeepCert ind28.target6.eps0.2 UNS 10.3 412.3 UNS 714.5 945.7
DeepCert ind28.target6.eps0.4 UNS 30.6 571.4 UNS 433.7 939.1
DeepCert ind28.target6.eps0.6 SAT 24.7 579.8 SAT 437.0 1892.1
DeepCert ind29.target6.eps0.2 UNS 10.7 408.6 UNS 597.4 782.6
DeepCert ind29.target6.eps0.4 SAT 24.9 571.6 SAT 687.2 1112.0
DeepCert ind29.target6.eps0.6 SAT 24.7 580.0 SAT 461.1 1248.9
DeepCert ind31.target6.eps0.2 UNS 25.6 551.3 UNS 549.4 949.7
DeepCert ind31.target6.eps0.4 SAT 24.2 565.1 SAT 309.5 1089.2
DeepCert ind31.target6.eps0.6 SAT 23.9 574.7 SAT 366.0 1268.6
DeepCert ind33.target6.eps0.2 UNS 24.0 550.0 UNS 804.9 1414.6
DeepCert ind33.target6.eps0.4 SAT 24.1 560.0 SAT 211.3 1409.0
DeepCert ind33.target6.eps0.6 SAT 24.0 573.9 SAT 357.8 757.1
DeepCert ind39.target6.eps0.2 UNS 9.9 406.6 UNS 300.2 603.1
DeepCert ind39.target6.eps0.4 UNS 30.6 559.5 UNS 1185.9 3033.1
DeepCert ind39.target6.eps0.6 UNS 33.0 573.1 UNS 1742.9 2882.3
DeepCert ind42.target6.eps0.2 UNS 10.7 406.5 UNS 603.8 940.5
DeepCert ind42.target6.eps0.4 UNS 39.0 563.5 UNS 1143.6 3472.4
DeepCert ind42.target6.eps0.6 UNS 53.1 572.3 UNS 1596.8 2047.7
DeepCert ind43.target6.eps0.2 UNS 10.7 407.9 UNS 340.0 602.7
DeepCert ind43.target6.eps0.4 UNS 29.7 565.4 UNS 1673.5 2244.1
DeepCert ind43.target6.eps0.6 UNS 33.3 574.9 UNS 1622.2 3670.7
DeepCert ind45.target6.eps0.2 UNS 10.6 406.9 UNS 428.6 781.9
DeepCert ind45.target6.eps0.4 UNS 31.6 563.4 UNS 1208.6 2048.1
DeepCert ind45.target6.eps0.6 UNS 25.5 572.8 UNS 2229.4 3658.6
DeepCert ind46.target6.eps0.2 UNS 9.8 404.5 UNS 1287.0 1890.3
DeepCert ind46.target6.eps0.4 UNS 26.2 559.8 UNS 1902.9 2328.9
DeepCert ind46.target6.eps0.6 UNS 40.0 565.2 UNS 951.6 2042.0
DeepCert ind47.target6.eps0.2 UNS 10.0 403.7 UNS 631.0 941.5
DeepCert ind47.target6.eps0.4 UNS 24.9 558.5 UNS 736.4 1258.9
DeepCert ind47.target6.eps0.6 UNS 31.8 568.1 UNS 1700.8 1771.6
DeepCert ind51.target6.eps0.2 UNS 10.2 404.7 UNS 365.8 1102.2
DeepCert ind51.target6.eps0.4 UNS 25.4 557.1 UNS 1807.4 2528.8
DeepCert ind51.target6.eps0.6 UNS 39.5 570.0 UNS 2099.4 1973.8
DeepCert ind53.target6.eps0.2 UNS 10.0 402.6 UNS 376.4 783.7
DeepCert ind53.target6.eps0.4 UNS 24.8 558.0 UNS 1429.3 2523.9
DeepCert ind53.target6.eps0.6 UNS 34.3 566.6 UNS 1480.0 1731.3
DeepCert ind55.target6.eps0.2 UNS 9.5 400.3 UNS 235.3 600.9
DeepCert ind55.target6.eps0.4 UNS 24.7 550.6 UNS 1388.7 1972.8
DeepCert ind55.target6.eps0.6 UNS 30.5 563.8 UNS 875.4 2053.1
DeepCert ind56.target6.eps0.2 UNS 10.1 401.2 UNS 579.3 1101.2
DeepCert ind56.target6.eps0.4 UNS 23.5 550.8 UNS 809.9 1137.7
DeepCert ind56.target6.eps0.6 UNS 38.9 564.6 UNS 1320.3 1728.9
DeepCert ind6.target6.eps0.2 UNS 10.8 410.2 UNS 417.5 1098.1
DeepCert ind6.target6.eps0.4 UNS 24.2 566.0 UNS 701.0 1576.8
DeepCert ind6.target6.eps0.6 SAT 27.4 574.4 SAT 734.7 1255.1
DeepCert ind60.target6.eps0.2 UNS 8.3 357.1 UNS 543.8 604.8
DeepCert ind60.target6.eps0.4 UNS 8.4 368.1 UNS 302.9 599.8
DeepCert ind60.target6.eps0.6 UNS 24.7 503.1 UNS 3366.6 2526.6
DeepCert ind66.target6.eps0.2 UNS 7.4 332.3 UNS 465.4 601.4
DeepCert ind66.target6.eps0.4 SAT 15.0 454.2 SAT 1079.6 1255.9
DeepCert ind66.target6.eps0.6 SAT 16.1 463.5 SAT 1648.3 2406.2
DeepCert ind67.target6.eps0.2 UNS 7.0 333.1 UNS 503.2 602.8
DeepCert ind67.target6.eps0.4 UNS 20.0 457.7 UNS 994.9 1110.3
DeepCert ind67.target6.eps0.6 SAT 17.1 466.9 SAT 2728.2 2375.9
DeepCert ind68.target6.eps0.2 UNS 6.2 304.1 UNS 722.2 607.0
DeepCert ind68.target6.eps0.4 UNS 6.6 313.9 UNS 679.7 949.3
DeepCert ind68.target6.eps0.6 UNS 16.1 423.4 UNS 2760.5 1776.7
DeepCert ind69.target6.eps0.2 UNS 6.3 318.5 UNS 583.5 604.9
DeepCert ind69.target6.eps0.4 UNS 14.5 437.6 UNS 657.7 948.4
DeepCert ind69.target6.eps0.6 UNS 16.6 446.0 UNS 3417.4 2102.0
DeepCert ind7.target6.eps0.2 UNS 10.1 411.1 UNS 394.2 781.0
DeepCert ind7.target6.eps0.4 UNS 26.0 569.1 UNS 1031.4 1734.9
DeepCert ind7.target6.eps0.6 SAT 27.2 573.7 SAT 1410.3 2172.1
DeepCert ind70.target6.eps0.2 UNS 6.4 321.3 UNS 721.9 604.4
DeepCert ind70.target6.eps0.4 UNS 12.9 438.6 UNS 756.1 787.1
DeepCert ind70.target6.eps0.6 UNS 18.3 447.7 UNS 2683.7 1617.1
DeepCert ind71.target6.eps0.2 UNS 6.8 325.3 UNS 731.5 607.8
DeepCert ind71.target6.eps0.4 UNS 7.3 340.8 UNS 1426.9 1421.8
DeepCert ind71.target6.eps0.6 UNS 17.9 463.3 UNS 744.2 941.4
DeepCert ind72.target6.eps0.2 UNS 7.6 336.0 UNS 536.4 604.2
DeepCert ind72.target6.eps0.4 UNS 7.6 351.9 UNS 1011.7 1107.8
DeepCert ind72.target6.eps0.6 UNS 18.1 479.6 UNS 934.7 944.2
DeepCert ind73.target6.eps0.2 UNS 7.9 356.9 UNS 683.5 605.1
DeepCert ind73.target6.eps0.4 UNS 8.2 369.1 UNS 826.2 942.2
DeepCert ind73.target6.eps0.6 UNS 18.8 504.3 UNS 1616.3 1736.5
DeepCert ind74.target6.eps0.2 UNS 8.0 361.6 UNS 722.7 606.0
DeepCert ind74.target6.eps0.4 UNS 8.3 370.2 UNS 433.7 781.5
DeepCert ind74.target6.eps0.6 UNS 19.9 507.2 UNS 1150.7 1734.0
DeepCert ind75.target6.eps0.2 UNS 7.8 352.6 UNS 802.0 943.3
DeepCert ind75.target6.eps0.4 UNS 7.7 363.2 UNS 423.9 601.3
DeepCert ind75.target6.eps0.6 UNS 19.5 495.6 UNS 482.7 780.5
DeepCert ind76.target6.eps0.2 UNS 7.7 346.7 UNS 705.4 604.7
DeepCert ind76.target6.eps0.4 UNS 7.5 359.8 UNS 1254.2 941.2
DeepCert ind76.target6.eps0.6 UNS 17.3 490.6 UNS 1929.3 1617.7
DeepCert ind77.target6.eps0.2 UNS 7.9 352.9 UNS 638.5 604.6
DeepCert ind77.target6.eps0.4 UNS 8.3 361.0 UNS 787.4 780.9
DeepCert ind77.target6.eps0.6 UNS 8.1 369.3 UNS 1240.1 1452.1
DeepCert ind78.target6.eps0.2 UNS 7.8 351.3 UNS 756.0 785.4
DeepCert ind78.target6.eps0.4 UNS 8.3 361.1 UNS 556.1 603.6
DeepCert ind78.target6.eps0.6 UNS 19.4 492.1 UNS 1791.8 2415.0
DeepCert ind79.target6.eps0.2 UNS 8.1 362.8 UNS 566.5 603.6
DeepCert ind79.target6.eps0.4 UNS 7.9 373.5 UNS 2526.8 1936.0
DeepCert ind79.target6.eps0.6 UNS 21.1 509.5 UNS 2616.3 1935.7
DeepCert ind8.target6.eps0.2 UNS 10.0 407.6 UNS 510.8 941.3
DeepCert ind8.target6.eps0.4 UNS 35.7 559.2 UNS 1365.8 1495.2
DeepCert ind8.target6.eps0.6 SAT 26.4 572.8 SAT 226.7 910.8
DeepCert ind80.target6.eps0.2 UNS 9.2 377.3 UNS 527.5 604.0
DeepCert ind80.target6.eps0.4 UNS 8.7 387.4 UNS 581.7 1257.7
DeepCert ind80.target6.eps0.6 UNS 33.3 530.1 UNS 1696.3 1892.8
DeepCert ind81.target6.eps0.2 UNS 8.5 370.9 UNS 355.7 602.4
DeepCert ind81.target6.eps0.4 UNS 9.0 379.6 UNS 1711.7 2054.9
DeepCert ind81.target6.eps0.6 UNS 27.0 519.2 UNS 3017.8 2612.1
DeepCert ind82.target6.eps0.2 UNS 8.2 366.6 UNS 939.1 944.8
DeepCert ind82.target6.eps0.4 UNS 8.4 375.6 UNS 848.6 1099.6
DeepCert ind82.target6.eps0.6 UNS 19.8 513.2 UNS 1232.6 1573.7
DeepCert ind83.target6.eps0.2 UNS 7.3 356.6 UNS 619.7 604.7
DeepCert ind83.target6.eps0.4 UNS 7.5 365.5 UNS 794.1 781.3
DeepCert ind83.target6.eps0.6 UNS 21.1 497.5 UNS 1028.2 1099.1
DeepCert ind84.target6.eps0.2 UNS 7.7 347.6 UNS 541.6 603.8
DeepCert ind84.target6.eps0.4 UNS 7.9 358.9 UNS 464.5 601.9
DeepCert ind84.target6.eps0.6 UNS 18.7 489.0 UNS 1290.3 1742.3
DeepCert ind9.target6.eps0.2 UNS 10.2 412.3 UNS 335.5 605.0
DeepCert ind9.target6.eps0.4 SAT 26.8 569.1 SAT 380.7 757.5
DeepCert ind9.target6.eps0.6 SAT 26.3 574.2 SAT 761.4 3307.1
NAP cls0.nap1 SAT 2.7 45.0 UNK 3601.0 1002.2
NAP cls0.nap10 UNK 3601.0 58.8 UNK 3601.0 1170.2
NAP cls0.nap11 UNK 3601.0 59.9 UNK 3601.0 1035.4
NAP cls0.nap12 SAT 2866.8 57.1 UNK 3601.0 1035.2
NAP cls0.nap13 UNK 3601.0 60.1 UNK 3601.0 1090.6
NAP cls0.nap14 SAT 2.6 45.2 UNK 3601.0 952.8
NAP cls0.nap15 UNK 3601.0 58.9 UNK 3601.0 1139.4
NAP cls0.nap16 UNK 3601.0 62.4 UNK 3601.0 1074.2
NAP cls0.nap17 UNK 3601.0 60.1 UNK 3601.1 1272.1
NAP cls0.nap18 UNK 3601.0 60.4 UNK 3601.1 1106.1
NAP cls0.nap19 UNK 3601.0 60.2 UNK 3601.0 1132.6
NAP cls0.nap2 UNK 3601.0 57.7 UNK 3601.0 1050.2
NAP cls0.nap20 UNK 3601.0 60.9 UNK 3601.1 1192.9
NAP cls0.nap21 SAT 4.0 48.2 UNK 3601.0 1197.1
NAP cls0.nap22 UNK 3601.0 58.6 UNK 3601.0 1024.7
NAP cls0.nap23 UNK 3601.0 57.7 UNK 3601.0 973.2
NAP cls0.nap24 UNK 3601.0 58.2 UNK 3601.1 1241.4
NAP cls0.nap25 UNK 3601.0 58.2 UNK 3601.0 1087.6
NAP cls0.nap26 SAT 276.0 56.5 UNK 3601.0 984.7
NAP cls0.nap27 UNK 3601.0 57.3 UNK 3601.0 1095.6
NAP cls0.nap28 UNK 3601.0 60.5 UNK 3601.0 1039.9
NAP cls0.nap29 SAT 21.8 58.8 UNK 3601.0 1034.3
NAP cls0.nap3 SAT 2.4 44.8 UNK 3601.0 1219.9
NAP cls0.nap30 SAT 170.6 56.2 UNK 3601.0 846.8
NAP cls0.nap31 UNK 3601.0 58.8 UNK 3601.0 1058.1
NAP cls0.nap32 UNK 3601.0 57.8 UNK 3601.1 1208.1
NAP cls0.nap33 UNK 3601.0 56.8 UNK 3601.0 753.6
NAP cls0.nap34 SAT 2387.7 58.2 UNK 3601.0 970.2
NAP cls0.nap35 SAT 1251.2 58.0 UNK 3601.0 1094.9
NAP cls0.nap36 UNK 3601.0 58.4 UNK 3601.1 1128.7
NAP cls0.nap37 UNK 3601.0 58.3 UNK 3601.0 1008.3
NAP cls0.nap38 UNK 3601.0 57.8 UNK 3601.0 988.0
NAP cls0.nap39 SAT 1132.7 58.7 UNK 3601.1 1298.8
NAP cls0.nap4 SAT 3.0 45.3 UNK 3601.0 1126.1
NAP cls0.nap40 UNK 3601.0 57.9 UNK 3601.0 1011.1
NAP cls0.nap41 UNK 3601.0 58.3 UNK 3601.0 1036.8
NAP cls0.nap42 UNK 3601.0 58.3 UNK 3601.1 1413.4
NAP cls0.nap43 UNK 3601.0 57.8 UNK 3601.1 1215.2
NAP cls0.nap44 SAT 122.8 57.8 UNK 3601.1 1372.7
NAP cls0.nap45 UNK 3601.0 58.9 UNK 3601.1 1435.9
NAP cls0.nap46 UNK 3601.0 58.0 UNK 3601.0 965.0
NAP cls0.nap47 UNK 3601.0 58.6 UNK 3601.1 1324.0
NAP cls0.nap48 UNK 3601.0 57.2 UNK 3601.0 1104.5
NAP cls0.nap49 UNK 3601.0 57.0 UNK 3601.0 928.7
NAP cls0.nap5 SAT 2.0 45.3 UNK 3601.1 1152.9
NAP cls0.nap50 UNK 3601.0 59.6 UNK 3601.1 1138.8
NAP cls0.nap51 UNK 3601.0 58.0 UNK 3601.0 998.7
NAP cls0.nap52 UNK 3601.0 58.9 UNK 3601.1 1334.7
NAP cls0.nap53 UNK 3601.0 58.5 UNK 3601.0 1115.1
NAP cls0.nap54 UNK 3601.0 58.6 UNK 3601.1 1257.3
NAP cls0.nap55 UNK 3601.0 61.5 UNK 3601.0 846.1
NAP cls0.nap56 UNK 3601.0 57.9 UNK 3601.1 1247.0
NAP cls0.nap57 UNK 3601.0 58.2 UNK 3601.0 1078.0
NAP cls0.nap58 UNK 3601.0 57.4 UNK 3601.0 1090.7
NAP cls0.nap59 UNK 3601.0 59.7 UNK 3601.0 947.0
NAP cls0.nap6 UNK 3601.0 60.7 UNK 3601.0 1051.6
NAP cls0.nap60 SAT 1194.9 58.0 UNK 3601.0 970.5
NAP cls0.nap61 UNK 3601.0 58.3 UNK 3601.0 1106.2
NAP cls0.nap62 UNK 3601.0 59.9 UNK 3601.0 912.8
NAP cls0.nap63 UNK 3601.0 57.9 UNK 3601.0 1073.0
NAP cls0.nap64 UNK 3601.0 60.2 UNK 3601.0 937.1
NAP cls0.nap65 UNK 3601.0 57.9 UNK 3601.0 726.5
NAP cls0.nap66 UNK 3601.0 60.5 UNK 3601.0 1064.6
NAP cls0.nap67 UNK 3601.0 57.1 UNK 3601.1 1391.0
NAP cls0.nap68 UNK 3601.0 58.5 UNK 3601.0 956.8
NAP cls0.nap69 UNK 3601.0 57.2 UNK 3601.1 1249.9
NAP cls0.nap7 SAT 2.5 45.1 UNK 3601.0 1106.7
NAP cls0.nap70 UNK 3601.0 57.9 UNK 3601.1 1110.0
NAP cls0.nap71 UNK 3601.0 59.0 UNK 3601.0 1027.2
NAP cls0.nap72 UNK 3601.0 57.8 UNK 3601.1 1166.1
NAP cls0.nap73 UNK 3601.0 58.9 UNK 3601.0 1160.1
NAP cls0.nap74 UNK 3601.0 56.7 UNK 3601.1 962.4
NAP cls0.nap75 UNK 3601.0 57.9 UNK 3601.0 1137.9
NAP cls0.nap76 UNK 3601.0 58.4 UNK 3601.1 1317.1
NAP cls0.nap8 UNK 3601.0 59.1 UNK 3601.0 1043.1
NAP cls0.nap9 SAT 1434.9 59.6 UNK 3601.0 1193.6
NAP cls1.nap77 UNK 3601.0 61.1 UNK 3601.1 1149.2
NAP cls1.nap78 UNK 3601.0 59.3 UNK 3601.0 1102.7
NAP cls1.nap79 SAT 18.0 58.3 UNK 3601.1 1157.8
NAP cls1.nap80 UNK 3601.0 61.2 UNK 3601.0 941.2
NAP cls1.nap81 UNK 3601.0 58.3 UNK 3601.0 981.8
NAP cls1.nap82 UNK 3601.0 62.9 UNK 3601.1 1297.8
NAP cls1.nap83 UNK 3601.0 60.4 UNK 3601.0 1002.1
NAP cls1.nap84 SAT 771.8 64.5 UNK 3601.0 1093.0
NAP cls1.nap85 UNK 3601.0 59.5 UNK 3601.0 1135.2
NAP cls1.nap86 SAT 6.8 55.0 UNK 3601.0 837.2
NAP cls1.nap87 UNK 3601.0 59.5 UNK 3601.1 1222.5
NAP cls1.nap88 UNK 3601.0 59.9 UNK 3601.0 1068.2
NAP cls2.nap100 UNK 3601.0 57.3 UNK 3601.1 1246.2
NAP cls2.nap101 SAT 1285.3 60.4 UNK 3601.0 1035.8
NAP cls2.nap102 UNK 3601.0 57.7 UNK 3601.0 1107.1
NAP cls2.nap103 UNK 3601.0 58.8 UNK 3601.0 1088.9
NAP cls2.nap104 UNK 3601.0 57.4 UNK 3601.1 1166.8
NAP cls2.nap105 UNK 3601.0 60.2 UNK 3601.0 1220.3
NAP cls2.nap89 UNK 3601.0 60.0 UNK 3601.0 1030.3
NAP cls2.nap90 UNK 3601.0 59.3 UNK 3601.1 1142.3
NAP cls2.nap91 UNK 3601.0 57.0 UNK 3601.1 1096.6
NAP cls2.nap92 UNK 3601.0 59.5 UNK 3601.1 1205.4
NAP cls2.nap93 UNK 3601.0 59.4 UNK 3601.0 1057.7
NAP cls2.nap94 UNK 3601.0 58.7 UNK 3601.0 634.0
NAP cls2.nap95 UNK 3601.0 58.5 UNK 3601.0 881.4
NAP cls2.nap96 UNK 3601.0 58.3 UNK 3601.0 1049.3
NAP cls2.nap97 UNK 3601.0 58.4 UNK 3601.1 1166.9
NAP cls2.nap98 UNK 3601.0 61.2 UNK 3601.0 912.4
NAP cls2.nap99 UNK 3601.0 62.4 UNK 3601.1 1257.3
NAP cls3.nap106 UNK 3601.0 60.6 UNK 3601.0 899.9
NAP cls3.nap107 SAT 378.4 62.7 UNK 3601.1 1267.5
NAP cls3.nap108 UNK 3601.0 64.8 UNK 3601.0 1093.3
NAP cls3.nap109 UNK 3601.0 58.6 UNK 3601.1 1428.3
NAP cls3.nap110 UNK 3601.0 64.0 UNK 3601.0 737.1
NAP cls3.nap111 UNK 3601.0 60.0 UNK 3601.0 1164.7
NAP cls3.nap112 SAT 2916.5 66.3 SAT 160.5 677.5
NAP cls3.nap113 UNK 3601.0 58.6 UNK 3601.0 1062.7
NAP cls3.nap114 UNK 3601.0 59.0 UNK 3601.1 1432.3
NAP cls3.nap115 SAT 3.1 47.2 UNK 3601.1 1046.5
NAP cls3.nap116 UNK 3601.0 60.4 UNK 3601.0 1010.6
NAP cls3.nap117 SAT 563.8 66.2 UNK 3601.0 931.4
NAP cls3.nap118 SAT 12.8 54.5 UNK 3601.0 926.1
NAP cls3.nap119 SAT 1497.3 63.6 UNK 3601.1 1147.0
NAP cls3.nap120 SAT 2320.7 59.9 SAT 3323.4 1127.4
NAP cls3.nap121 UNK 3601.0 58.6 UNK 3601.0 802.9
NAP cls3.nap122 UNK 3601.0 63.0 UNK 3601.0 789.5
NAP cls3.nap123 SAT 529.6 62.8 UNK 3601.0 1127.5
NAP cls3.nap124 UNK 3601.0 62.9 UNK 3601.1 1268.4
NAP cls3.nap125 SAT 2.6 45.0 UNK 3601.1 1013.2
NAP cls3.nap126 SAT 824.0 61.1 UNK 3601.1 1405.9
NAP cls3.nap127 UNK 3601.0 58.2 UNK 3601.1 1350.5
NAP cls3.nap128 UNK 3601.0 61.9 UNK 3601.1 1251.3
NAP cls3.nap129 UNK 3601.0 60.5 UNK 3601.0 1064.2
NAP cls3.nap130 UNK 3601.0 58.7 UNK 3601.0 886.6
NAP cls3.nap131 UNK 3601.0 65.8 UNK 3601.0 1146.7
NAP cls3.nap132 SAT 2.0 45.2 SAT 2882.8 955.8
NAP cls3.nap133 UNK 3601.0 62.4 UNK 3601.0 949.1
NAP cls3.nap134 UNK 3601.0 60.9 UNK 3601.1 1016.9
NAP cls3.nap135 UNK 3601.0 63.6 UNK 3601.0 1067.0
NAP cls3.nap136 UNK 3601.0 63.2 UNK 3601.0 928.2
NAP cls3.nap137 UNK 3601.0 59.4 UNK 3601.0 1060.5
NAP cls3.nap138 UNK 3601.0 57.9 UNK 3601.1 1055.6
NAP cls3.nap139 UNK 3601.0 59.0 UNK 3601.0 940.7
NAP cls3.nap140 UNK 3601.0 61.7 SAT 2539.1 1029.8
NAP cls3.nap141 UNK 3601.0 59.1 UNK 3601.0 1010.3
NAP cls3.nap142 SAT 2442.3 63.1 UNK 3601.1 1276.4
NAP cls3.nap143 UNK 3601.0 58.8 UNK 3601.1 1107.5
NAP cls3.nap144 UNK 3601.0 58.5 UNK 3601.0 980.8
NAP cls3.nap145 UNK 3601.0 59.4 SAT 81.3 905.3
NAP cls3.nap146 UNK 3601.0 65.0 SAT 842.4 933.9
NAP cls3.nap147 UNK 3601.0 58.3 UNK 3601.1 1170.6
NAP cls3.nap148 UNK 3601.0 62.2 UNK 3601.0 1140.8
NAP cls3.nap149 SAT 2.9 44.9 UNK 3601.0 978.0
NAP cls3.nap150 UNK 3601.0 57.6 UNK 3601.1 1595.4
NAP cls3.nap151 UNK 3601.0 59.3 UNK 3601.0 896.0
NAP cls3.nap152 UNK 3601.0 59.0 UNK 3601.0 1099.5
NAP cls3.nap153 UNK 3601.0 66.4 UNK 3601.0 995.1
NAP cls3.nap154 UNK 3601.0 58.3 UNK 3601.0 1167.0
NAP cls3.nap155 UNK 3601.0 58.2 UNK 3601.0 1098.2
NAP cls3.nap156 SAT 3.7 46.0 UNK 3601.0 1009.3
NAP cls3.nap157 UNK 3601.0 57.1 UNK 3601.0 1106.1
NAP cls3.nap158 UNK 3601.0 56.5 SAT 49.0 625.2
NAP cls3.nap159 UNK 3601.0 59.6 UNK 3601.0 1099.7
NAP cls3.nap160 UNK 3601.0 64.6 UNK 3601.0 963.7
NAP cls3.nap161 UNK 3601.0 63.6 UNK 3601.1 950.2
NAP cls3.nap162 UNK 3601.0 58.8 UNK 3601.0 1014.4
NAP cls3.nap163 UNK 3601.0 59.2 UNK 3601.0 1058.5
NAP cls3.nap164 UNK 3601.0 60.8 UNK 3601.0 1028.0
NAP cls3.nap165 UNK 3601.0 58.3 UNK 3601.1 1209.0
NAP cls3.nap166 UNK 3601.0 58.7 UNK 3601.1 1132.5
NAP cls3.nap167 UNK 3601.0 58.6 UNK 3601.0 962.5
NAP cls3.nap168 UNK 3601.0 63.2 UNK 3601.1 1217.6
NAP cls3.nap169 SAT 788.3 60.8 UNK 3601.1 1200.4
NAP cls3.nap170 UNK 3601.0 62.2 UNK 3601.0 820.0
NAP cls3.nap171 UNK 3601.0 59.5 UNK 3601.1 1468.8
NAP cls3.nap172 UNK 3601.0 60.0 UNK 3601.0 993.7
NAP cls4.nap173 UNK 3601.0 58.1 UNK 3601.0 907.6
NAP cls4.nap174 UNK 3601.0 60.1 UNK 3601.0 975.6
NAP cls4.nap175 UNK 3601.0 60.4 UNK 3601.1 1343.9
NAP cls4.nap176 UNK 3601.0 62.6 UNK 3601.0 1095.5
NAP cls4.nap177 UNK 3601.0 57.2 UNK 3601.1 1105.3
NAP cls4.nap178 UNK 3601.0 59.2 UNK 3601.0 982.6
NAP cls4.nap179 UNK 3601.0 58.4 UNK 3601.0 1010.6
NAP cls4.nap180 SAT 1340.1 67.8 UNK 3601.1 1028.0
NAP cls4.nap181 UNK 3601.0 56.2 UNK 3601.1 1180.4
NAP cls4.nap182 UNK 3601.0 58.1 SAT 171.3 530.2
NAP cls4.nap183 UNK 3601.0 60.9 UNK 3601.0 890.4
NAP cls4.nap184 UNK 3601.0 64.5 UNK 3601.0 968.7
NAP cls4.nap185 SAT 1143.3 65.3 UNK 3601.1 1083.6
NAP cls4.nap186 SAT 1716.8 58.7 UNK 3601.0 1089.6
NAP cls4.nap187 UNK 3601.0 58.7 UNK 3601.1 1143.1
NAP cls4.nap188 UNK 3601.0 59.2 UNK 3601.1 1324.8
NAP cls4.nap189 UNK 3601.0 59.0 UNK 3601.1 1257.3
NAP cls4.nap190 UNK 3601.0 57.7 UNK 3601.0 1255.5
NAP cls4.nap191 UNK 3601.0 58.1 UNK 3601.1 1081.5
NAP cls4.nap192 UNK 3601.0 57.3 UNK 3601.0 986.5
NAP cls4.nap193 UNK 3601.0 58.1 UNK 3601.1 1240.3
NAP cls4.nap194 UNK 3601.0 58.4 UNK 3601.0 824.6
NAP cls4.nap195 SAT 1.8 45.3 UNK 3601.1 1105.0
NAP cls4.nap196 UNK 3601.0 59.0 UNK 3601.0 1081.9
NAP cls4.nap197 UNK 3601.0 55.7 UNK 3601.1 1350.3
NAP cls4.nap198 SAT 2374.8 62.2 UNK 3601.1 1241.2
NAP cls4.nap199 UNK 3601.0 57.5 UNK 3601.0 793.0
NAP cls4.nap200 UNK 3601.0 59.2 UNK 3601.0 1015.7
NAP cls4.nap201 UNK 3601.0 58.7 UNK 3601.0 997.9
NAP cls4.nap202 UNK 3601.0 60.1 UNK 3601.0 812.3
NAP cls4.nap203 UNK 3601.0 60.0 UNK 3601.1 1084.9
NAP cls4.nap204 UNK 3601.0 58.4 UNK 3601.1 1514.7
NAP cls4.nap205 UNK 3601.0 58.8 UNK 3601.0 967.2
NAP cls4.nap206 UNK 3601.0 58.2 UNK 3601.0 949.8
NAP cls4.nap207 UNK 3601.0 55.7 UNK 3601.1 1294.3
NAP cls4.nap208 UNK 3601.0 61.0 UNK 3601.1 961.4
NAP cls4.nap209 UNK 3601.0 58.5 UNK 3601.1 1251.7
NAP cls4.nap210 UNK 3601.0 56.3 UNK 3601.0 774.3
NAP cls4.nap211 UNK 3601.0 58.1 UNK 3601.0 992.3
NAP cls4.nap212 UNK 3601.0 58.0 UNK 3601.1 1356.3
NAP cls4.nap213 UNK 3601.0 60.5 UNK 3601.0 1095.2
NAP cls4.nap214 UNK 3601.0 58.1 UNK 3601.1 1091.5
NAP cls4.nap215 UNK 3601.0 58.9 UNK 3601.1 1154.9
NAP cls4.nap216 UNK 3601.0 58.9 UNK 3601.1 1044.2
NAP cls4.nap217 UNK 3601.0 57.6 UNK 3601.0 887.3
NAP cls4.nap218 UNK 3601.0 58.8 UNK 3601.1 1367.5
NAP cls4.nap219 UNK 3601.0 60.0 UNK 3601.1 1085.1
NAP cls4.nap220 UNK 3601.0 57.9 UNK 3601.1 1167.6
NAP cls4.nap221 UNK 3601.0 57.4 UNK 3601.0 1051.8
NAP cls4.nap222 UNK 3601.0 60.5 UNK 3601.1 1407.6
NAP cls4.nap223 SAT 2504.9 66.3 UNK 3601.0 1072.7
NAP cls4.nap224 UNK 3601.0 62.6 UNK 3601.1 1041.2
NAP cls4.nap225 UNK 3601.0 57.1 UNK 3601.1 1136.6
NAP cls4.nap226 UNK 3601.0 57.3 UNK 3601.1 1185.4
NAP cls4.nap227 UNK 3601.0 60.4 UNK 3601.1 1159.9
NAP cls4.nap228 UNK 3601.0 57.9 UNK 3601.1 1199.0
NAP cls4.nap229 UNK 3601.0 57.4 UNK 3601.1 1493.2
NAP cls4.nap230 UNK 3601.0 58.0 UNK 3601.0 937.2
NAP cls4.nap231 UNK 3601.0 56.0 UNK 3601.0 822.6
NAP cls4.nap232 UNK 3601.0 58.6 UNK 3601.1 1199.7
NAP cls4.nap233 UNK 3601.0 57.9 UNK 3601.0 973.0
NAP cls4.nap234 UNK 3601.0 57.7 SAT 15.6 265.8
NAP cls4.nap235 UNK 3601.0 59.7 UNK 3601.1 1155.8
VeriX index0 OK 661.1 923.8 OK 1483.1 1048.9
VeriX index1 OK 655.7 924.8 OK 1504.1 1045.5
VeriX index2 OK 1316.9 907.5 OK 3257.5 1045.6
VeriX index3 OK 1046.0 907.4 OK 2748.7 1045.5
VeriX index4 OK 1749.5 693.3 TO 3601.0 502.1
VeriX index5 OK 1258.3 906.5 OK 3101.4 1047.2
VeriX index6 OK 749.0 692.8 OK 2558.8 1041.0
VeriX index7 OK 752.8 906.1 OK 1854.0 509.2
VeriX index9 OK 892.4 692.8 OK 2213.8 502.1
VeriX index10 OK 774.3 709.3 OK 1665.6 501.9
VeriX index11 OK 1586.7 693.2 TO 3601.0 502.2
VeriX index12 OK 864.9 693.2 OK 2642.2 1048.7
VeriX index13 OK 1958.5 692.0 TO 3601.0 502.0
VeriX index14 OK 758.8 906.6 OK 1556.6 503.3
VeriX index15 OK 720.3 709.9 OK 1791.4 502.0
VeriX index16 OK 1364.6 709.1 OK 2818.4 501.8
VeriX index17 OK 615.7 707.9 OK 1449.7 502.3
VeriX index19 OK 2473.9 708.3 TO 3601.0 502.1
VeriX index20 OK 762.2 710.2 OK 2091.4 1050.2
VeriX index21 OK 863.7 906.0 OK 2303.0 502.5
VeriX index22 OK 3370.4 692.0 TO 3601.0 503.4
VeriX index23 OK 760.6 709.6 OK 1715.9 502.1
VeriX index24 OK 2586.4 693.5 TO 3601.0 501.9
VeriX index25 OK 1838.4 708.1 OK 2197.4 502.5
VeriX index26 OK 857.7 692.3 OK 2392.8 502.4
VeriX index27 OK 1119.1 709.7 TO 3601.0 501.8
VeriX index28 OK 745.0 692.8 OK 1961.3 501.8
VeriX index29 OK 725.4 691.6 OK 1484.9 501.8
VeriX index30 OK 767.2 692.9 OK 1703.9 502.7
VeriX index31 OK 1300.5 708.3 OK 1955.8 501.9
VeriX index32 OK 700.7 691.5 OK 1627.3 501.8
VeriX index34 OK 673.7 710.0 OK 1409.0 502.1
VeriX index35 OK 919.1 692.6 OK 1722.8 501.6
VeriX index36 OK 936.9 691.9 OK 1508.8 501.2
VeriX index37 OK 1318.9 692.3 OK 1942.0 502.3
VeriX index38 OK 499.3 709.6 OK 1100.8 502.4
VeriX index39 OK 1249.6 691.7 OK 2102.3 502.3
VeriX index40 TO 3601.1 710.2 TO 3601.0 501.7
VeriX index41 OK 1445.9 693.1 OK 1641.0 502.0
VeriX index42 OK 851.1 692.6 OK 2096.1 502.0
VeriX index43 OK 1336.0 710.1 OK 3025.6 502.0
VeriX index44 OK 616.4 709.4 OK 1329.1 501.7
VeriX index45 OK 795.2 692.6 OK 1860.3 502.2
VeriX index46 OK 620.2 709.9 OK 1274.5 501.6
VeriX index47 OK 724.8 694.3 OK 1715.8 502.6
VeriX index48 OK 737.1 707.9 OK 1684.1 502.9
VeriX index49 OK 3497.3 708.8 TO 3601.0 501.6
VeriX index50 OK 1277.3 709.7 OK 2965.5 501.6
VeriX index51 OK 798.4 692.8 OK 1804.2 502.2
VeriX index52 OK 830.5 708.6 OK 1957.1 502.0
VeriX index53 OK 1753.4 709.8 TO 3601.0 501.6
VeriX index54 OK 747.4 708.7 OK 1567.8 502.6
VeriX index55 OK 755.0 692.1 OK 1829.7 501.7
VeriX index56 OK 1157.0 691.6 TO 3600.0 501.5
VeriX index57 OK 1514.6 709.0 OK 2715.4 501.6
VeriX index58 OK 823.9 692.3 OK 2200.0 502.4
VeriX index59 OK 3572.4 693.4 TO 3601.0 501.6
VeriX index60 OK 671.4 691.8 OK 1638.1 501.5
VeriX index61 OK 504.7 708.7 OK 1103.4 503.1
VeriX index62 OK 562.4 691.7 OK 1212.7 502.2
VeriX index64 OK 676.9 691.6 OK 1476.5 501.7
VeriX index65 OK 716.3 691.9 OK 2251.1 502.0
VeriX index66 OK 682.0 691.4 OK 2301.6 501.0
VeriX index67 OK 1602.9 707.8 TO 3601.0 501.8
VeriX index68 OK 759.6 692.6 OK 1825.2 502.1
VeriX index69 OK 734.2 691.5 OK 1700.0 502.3
VeriX index70 OK 1365.4 692.7 OK 1947.7 502.8
VeriX index71 OK 845.3 693.2 OK 1972.6 502.4
VeriX index72 OK 758.1 692.5 OK 1701.5 501.1
VeriX index73 OK 993.0 692.5 OK 2224.3 502.1
VeriX index74 OK 1251.6 708.8 OK 1902.1 501.9
VeriX index75 OK 927.2 708.1 OK 1841.3 501.3
VeriX index76 OK 745.1 708.2 OK 1898.4 501.2
VeriX index77 OK 3492.8 692.3 TO 3601.0 502.6
VeriX index78 OK 744.9 707.5 OK 1777.7 501.9
VeriX index79 OK 688.8 692.6 OK 1527.9 502.3
VeriX index80 OK 717.6 708.5 OK 1633.8 502.4
VeriX index81 OK 856.9 708.0 OK 2279.7 502.1
VeriX index82 OK 739.9 692.4 OK 1741.3 502.1
VeriX index83 OK 811.2 708.6 OK 1831.7 501.8
VeriX index84 OK 1113.8 710.0 OK 3061.4 502.0
VeriX index85 OK 750.3 691.9 OK 1853.2 502.1
VeriX index86 OK 999.6 692.1 OK 1896.5 501.7
VeriX index87 OK 576.1 709.2 OK 1265.4 502.9
VeriX index88 OK 1439.6 692.8 OK 2453.0 501.4
VeriX index89 OK 1196.4 692.9 OK 2290.9 503.0
VeriX index90 OK 740.4 708.8 OK 1842.2 502.9
VeriX index91 OK 863.8 693.4 OK 2204.0 501.3
VeriX index92 OK 680.2 692.3 OK 1653.2 501.5
VeriX index93 OK 707.9 692.1 OK 1608.4 502.3
VeriX index94 OK 778.3 708.8 OK 1637.0 502.8
VeriX index95 OK 710.4 707.5 OK 1863.5 503.2
VeriX index96 OK 1334.6 707.9 OK 1689.1 502.4
VeriX index97 OK 445.2 709.5 OK 979.6 501.7
VeriX index98 OK 683.9 708.5 OK 1500.1 501.9
VeriX index99 OK 1392.8 693.0 OK 3130.3 501.8
VeriX index100 OK 1505.1 708.6 TO 3601.0 502.0
VeriX index101 TO 3601.1 692.4 TO 3601.0 503.0
VeriX index102 OK 735.2 708.9 OK 1683.7 502.6
VeriX index103 OK 1107.1 692.4 TO 3601.0 501.8