### v2

Abstract
This paper serves as a comprehensive system description of version 2.0 of the Marabou framework for formal analysis of neural networks. We discuss the tool’s architectural design and highlight the major features and components introduced since its initial release.
1 Introduction
With the increasing pervasiveness of deep neural networks (DNNs), the formal analysis of DNNs has become a burgeoning research field within the formal methods community. Multiple DNN reasoners have been proposed in the past few years, including $\alpha$-$\beta$-CROWN [79, 67, 75], ERAN [56, 57, 58], Marabou [40], MN-BaB [18], NNV [62, 44], nnenum [4], VeriNet [30, 29], and many others.
We focus here on the Marabou [40] tool, which has been used by the research community in a wide range of formal DNN reasoning applications (e.g., [26, 12, 46, 19, 60, 74, 9, 20, 65, 76, 32, 43], inter alia). Initially, Marabou was introduced as a from-scratch re-implementation of the Reluplex [39] decision procedure, with a native linear programming engine and limited support for DNN-level reasoning. Over the years, fundamental changes have been made to the tool, not only from an algorithmic perspective but also to its engineering and implementation.
This paper introduces version 2.0 of Marabou. Compared to its predecessor, Marabou 2.0: (i) employs a new build/test system; (ii) has an optimized core system architecture; (iii) runs an improved decision procedure and abstract interpretation techniques; (iv) handles a wider range of activation functions; (v) supports proof production; (vi) supports additional input formats; and (vii) contains a more powerful Python API. Due to these changes, the original system description [40] no longer gives an accurate account of the tool’s current capabilities. Our goal in this paper is to close this gap and provide a comprehensive description of the current Marabou system. We highlight the major features introduced since the initial version, describe a few of its many recent uses, and report on its performance, as demonstrated by the VNN-COMP’23 results and additional runtime comparisons against an early version of Marabou.
2 Architecture and Core Components
C++ APICLIPython APICEGAR SolverMulti-thread ManagerAbstract InterpretationsSimplex EngineSMT Solver(MI)LP InterfaceLinear Solving EngineSMT CoreContext-dependentData StructuresNetwork-level ReasonerPreprocessorProof ModuleBack EndEngineNon-linear Constraints(In)equationsPiecewise-linear ConstraintsBoundsInput QueryFront End
Figure 1: High-level overview of Marabou 2.0’s system architecture.
In this section, we discuss the core components of Marabou 2.0. An overview of its system architecture is given in Figure 1. At a high level, Marabou performs satisfiability checking on a set of linear and non-linear constraints, supplied through one of the front-end interfaces. The constraints typically represent a verification query over a neural network and are stored in an InputQuery object. We distinguish variable bounds from other linear constraints, and piecewise-linear constraints (which can be reduced to linear constraints via case analysis) from more general, non-linear constraints.
Variables are represented as consecutive indices starting from 0. (In)equations are represented as Equation objects. Piecewise-linear constraints are represented by objects of classes that inherit from the PiecewiseLinearConstraint abstract class. The abstract class defines the key interface methods that are implemented in each sub-class. This way, all piecewise-linear constraints are handled uniformly in the back end. Similarly, each other type of non-linear constraint is implemented as a sub-class of the new NonlinearConstraint abstract class. Initially, Marabou only supported the ReLU and Max constraints. In Marabou 2.0, over ten types of non-linear constraints (listed in Appendix 0.A.1) are supported.
2.1 Engine
The centerpiece of Marabou is called the Engine, which reasons about the satisfiability of the input query. The engine consists of several components: the Preprocessor, which performs rewrites and simplifications; the Network-level Reasoner, which maintains the network architecture and performs all analyses that require this knowledge; the SMT Solver, which houses complete decision procedures for sets of linear and piecewise-linear constraints; and the (MI)LP Interface, which manages interactions with external (MI)LP solvers for certain optional solving modes as explained below.
Two additional modules are built on top of the Engine. The Multi-thread Manager spawns multiple Engine instances to take advantage of multiple processors. The CEGAR Solver performs incremental linearization [72, 13] for non-linear constraints that cannot be precisely handled by the SMT Solver.
2.1.1 Preprocessor.
Every verification query first goes through multiple preprocessing passes, which normalize, simplify, and rewrite the query. One new normalizing pass introduces auxiliary variables and entailed linear constraints for each of the piecewise-linear constraints, so that case splits on the piecewise-linear constraints can be represented as bound updates and consequently do not require adding new equations. [^1] For example, for a piece-wise linear constraint $y=\max(x_{1},x_{2})$, we would introduce $c_{1}:y-x_{1}=a_{1}\land a_{1}\geq 0\land y-x_{2}=a_{2}\land a_{2}\geq 0$, where $a_{1}$ and $a_{2}$ are fresh variables. This way, case splits on this constraint can be represented as $c_{2}:a_{1}\leq 0$ and $c_{3}:a_{2}\leq 0$, respectively. This preprocessing pass preserves satisfiability because the original constraint is equisatisfiable to $c_{1}\land(c_{2}\lor c_{3})$. This accelerates the underlying Simplex engine, as explained in the SMT Solver section below. Another significant preprocessing pass involves iterative bound propagation over all constraints. In this process, piecewise linear constraints might collapse into linear constraints and be removed. This pass was present in Marabou 1.0, but could become a runtime bottleneck; whereas Marabou 2.0 employs a data structure optimization that leads to a $\sim$60x speed up. Finally, the preprocessor merges any variables discovered to be equal to each other and also eliminates any constant variables. This results in updates to the variable indices, and therefore a mapping from old indices to new ones needs to be maintained for retrieving satisfying assignments.
2.1.2 SMT Solver.
The SMT Solver module implements a sound and complete, lazy-DPLL(T)-based procedure for deciding the satisfiability of a set of linear and piecewise-linear constraints. It performs case analysis on the piecewise-linear constraints and, at each search state, employs a specialized procedure to iteratively search for an assignment satisfying both the linear and non-linear constraints.
Presently, the DeepSoI procedure [69] has replaced the Reluplex procedure [39, 40] as Marabou’s default procedure to run at each search state. The former provably converges to a satisfying assignment (if it exists) and empirically consistently outperforms the latter. DeepSoI extends the canonical sum-of-infeasibilities method in convex optimization [10], which determines the satisfiability of a set of linear constraints by minimizing a cost function that represents the total violation of the constraints by the current assignment. The constraints are satisfiable if and only if the optimal value is 0. Similarly, DeepSoI formulates a cost function that represents the total violation of the current piecewise-linear constraints and uses a convex solver to stochastically minimize the cost function with respect to the convex relaxation of the current constraints. In addition, DeepSoI also informs the branching heuristics of the SMT Core, which performs a case split on the piecewise-linear constraint with the largest impact (measured by the pseudocost metric [69]) on the cost function. The DeepSoI procedure is implemented for all supported piecewise-linear activation functions. The convex solver can be instantiated either with the native Simplex engine or with an external LP solver via the (MI)LP interface (detailed below). The latter can be more efficient but requires the use of external commercial solvers.
One optimization in Marabou 2.0’s Simplex engine is that once the tableau has been initialized, it avoids introducing any new equations — a costly operation that requires re-computing the tableau from scratch. This is achieved by implementing case-splitting and backtracking as updates on variable bounds (as mentioned above), which only requires minimal updates to the tableau state. By our measure, this optimization reduces the runtime of the Simplex engine by over 50%. Moreover, the memory footprint of the solver is also drastically decreased, as the SMT Core no longer needs to save the entire tableau state during case-splitting (to be restored during backtracking).
2.1.3 Network-level Reasoner.
Over the past few years, numerous papers (e.g.,  [66, 57, 80, 50, 78], inter alia) have proposed abstract interpretation techniques that rely on network-level reasoning (e.g., propagating the input bounds layer by layer to tighten output bounds). These analyses can be viewed as a stand-alone, incomplete DNN verification procedure, or as in-processing bound tightening passes for the SMT Solver. Marabou 2.0 features a brand new NetworkLevelReasoner class that supports this type of analysis. The class maintains the neural network topology as a directed acyclic graph, where each node is a Layer object. The Layer class records key information such as weights, source layers, and mappings between neuron indices and variable indices. Currently, seven different analyses are implemented: [i] 1. interval bound propagation [24]; 2. symbolic bound propagation [66]; 3. DeepPoly/CROWN analysis [57, 80]; 4. LP-based bound tightening [61]; 5. Forward-backward analysis [70]; 6. MILP-based bound tightening [61]; and 7. iterative propagation [68]. Analyses 2–7 are implemented in a parallelizable manner, and analyses 4–7 require calls to an external LP solver. By default, the DeepPoly/CROWN analysis is performed. The Network-level Reasoner tightly interleaves with the SMT Solver: the network-level reasoning is executed any time a new search state is reached (with the most up-to-date variable bounds), and the derived bound tightenings are immediately fed back to the search procedure.
It is noteworthy that the user does not have to explicitly provide the neural network topology to enable network-level reasoning. Instead, the network architecture is automatically inferred from the given set of linear and non-linear constraints, via the constructNetworkLevelReasoner method in the InputQuery class. The Network-level Reasoner is only initialized if such inference is successful. Apart from the abstract interpretation passes, the Network-level Reasoner can also evaluate concrete inputs. This is used to implement the LP-based bound tightening optimization introduced by the NNV tool [62].
2.1.4 (MI)LP Interface.
Marabou can now optionally be configured to invoke the Gurobi Optimizer [27], a state-of-the-art Mixed Integer Linear Programming (MILP) solver. The GurobiWrapper class contains methods to construct a MILP problem and invoke the solver. The MILPEncoder class is in charge of encoding the current set of linear and non-linear constraints as (MI)LP constraints. Piecewise-linear constraints can either be encoded precisely, or replaced with a convex relaxation, resulting in a linear program. For other non-linear constraints, only the latter option is available. The (MI)LP interface presently has three usages in the code base. Two have already been mentioned, i.e., in some of the abstract interpretation passes and optionally in the DeepSoI procedure. Additionally, when Marabou is compiled with Gurobi, a --milp mode is available, in which the Engine performs preprocessing and abstract interpretation passes, and then directly encodes the verification problem as a MILP problem to be solved by Gurobi. The mode is motivated by the observation that the performance of Gurobi and the SMT Solver can be complementary [69, 59].
2.1.5 Multi-thread Manager.
Parallelization is an important way to improve verification efficiency. Marabou supports two modes of parallelization, both managed by the new MultiThreadManager class: the split-and-conquer mode [68] and the portfolio mode. In the split-and-conquer mode, the original query is dynamically partitioned and re-partitioned into independent sub-queries, to be handled by idle workers. The partitioning strategy is implemented as a sub-class of the QueryDivider abstract class. Currently, two strategies are available: one partitions the intervals of the input variables; the other splits on piecewise linear constraints. By default, the former is used only when the input dimension is less than or equal to ten. In the portfolio mode, each worker solves the same query with a different random seed, which takes advantage of the stochastic nature of the DeepSoI procedure. Developing an interface to define richer kinds of portfolios is work in progress.
2.1.6 CEGAR Solver.
While the DNN verification community has by and large focused on piecewise-linear activation functions, other classes of non-linear connections exist and are commonly used for certain architectures [33, 64]. Apart from introducing support for non-linear constraints in the Preprocessor and the Network-level Reasoner, the latest Marabou version also incorporates a counter-example guided abstraction refinement (CEGAR) solving mode [72], based on incremental linearization [13] to enable more precise reasoning about non-linear constraints that are not piecewise linear. Currently, the CEGAR solver only supports Sigmoid and Tanh, but the module can be extended to handle other activation functions.
2.2 Context-Dependent Data-Structures
When performing a case split or backtracking to a previous search state, the SMT Core needs to save or restore information such as variable bounds and the phase status of each piecewise-linear constraint (e.g., is a ReLU currently active, inactive, or unfixed). To efficiently support these operations, Marabou 2.0 uses the notion of a context level (borrowed from the CVC4 SMT solver [6]), and stores the aforementioned information in context-dependent data structures. These data structures behave similarly to their standard counterparts, except that they are associated with a context level and automatically save and restore their state as the context increases or decreases. This major refactoring has greatly simplified the implementation of saving and restoring solver states and is an important milestone in an ongoing effort to integrate a full-blown Conflict-Driven Clause-Learning (CDCL) mechanism into Marabou.
2.3 Proof Module
A proof module has recently been introduced into Marabou, enabling it to optionally produce proof certificates after an unsatisfiable (UNSAT) [36] result. This is common practice in the SAT and SMT communities and is aimed at ensuring solver reliability. Marabou produces proof certificates based on a constructive variant of the Farkas lemma [63], which ensures the existence of a proof vector that witnesses the unsatisfiability of a linear program. Specifically, the proof vector corresponds to a linear equation that is violated by the variable bounds [36]. The full certificate of UNSAT is comprised of a proof tree, whose nodes represent the search states explored during the solving. Each node may contain a list of lemmas that are used as additional constraints in its descendent nodes; and each leaf node contains the proof vector for the unsatisfiability of the corresponding sub-query. The lemmas encapsulate some of the variable bounds, newly derived by the piecewiese-linear constraints of the query, and require their own witnesses (i.e., proof vectors). The BoundExplainer class is responsible for constructing all proof vectors, for updating them during execution, and for appending them to the node. The proof tree itself is implemented using the UnsatCertificateNode class.
When the solver is run in proof-production mode, the Proof module closely tracks the steps of the SMT Solver module and constructs the proof tree on the fly: new nodes are added to the tree whenever a case split is performed; and a new proof vector is generated whenever a lemma is learned or UNSAT is derived for a sub-query. If the Engine concludes that the entire query is UNSAT, a proof checker (implemented as an instance of the Checker class) will be triggered to certify the proof tree. It does so by traversing the tree and certifying the correctness of the lemmas and the unsatisfiability of the leaf nodes. A formally verified and precise proof-checker is currently under development [16]. Note that, currently, proof production mode is only compatible with a subset of the features supported by Marabou. Adding support for the remaining features (e.g., for the parallel solving mode) is an ongoing endeavor.
2.4 Front End
Marabou provides interfaces to prepare input queries and invoke the back-end solver in multiple ways. The Marabou executable can be run on the command line, taking in network/property/query files in supported formats. The Python and C++ APIs support this functionality as well, but also contain methods to add arbitrary linear and (supported) non-linear constraints. In addition, a layer on top of the Python API was added to Marabou 2.0 which allows users to define constraints in a more Pythonic manner, resulting in more succinct code. For example, suppose one wants to check whether the first output of a network (stored in the ONNX format) can be less than or equal to half of its second output, when the first input is greater than or equal to $0.1$. Figure 2(a) shows how to perform this check with the base Python API, while Figure 2(b) exhibits the “Pythonic” API.
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.setLowerBound(X[0], 0.1)
Q.addInequality([Y[0], Y[1]], [1, -0.5], 0)
Q.solve()
((a)) The base Python API
Q = Marabou.read_onnx("model.onnx")
X, Y = Q.inputVars[0], Q.outputVars[0]
Q.addConstraint(Var(X[0]) >= 0.1)
Q.addConstraint(Var(Y[0]) <= 0.5 * Var(Y[1]))
Q.solve()
((b)) The “Pythonic” API
Figure 2: Two ways to define the same verification query through the Python API.
Typically, a query consists of the encoding of (one or several) neural networks and the encoding of a property on the network(s). To encode a neural network, the user has two options: 1) pass in a neural network file to be parsed by one of the neural network parsers; or 2) manually add constraints to encode the neural network. The main network format for Marabou 2.0 is now ONNX, towards which the neural network verification community is converging. The NNet format and the Tensorflow protobuf format are still supported but will likely be phased out in the long run. To encode the property on top of the neural network encoding, the user can 1) pass in a property file to be parsed by one of the property parsers; or 2) manually encode the property. Currently Marabou has two property parsers, one for a native property file format [40], and a new one for the VNN-LIB format, supporting the standardization effort of the community.
In addition to the aforementioned network and property file formats, Marabou also supports a native query file format that describes a set of linear and non-linear constraints. This can be dumped/parsed from all interfaces.
2.5 Availability, License, and Installation
Marabou is available under the permissive modified BSD open-source license, and runs on Linux and macOS machines. The tool can be built from scratch using CMake. Marabou is now also available on The Python Package Index (PyPI) and can be installed through pip. The latest version of Marabou is available at: https://github.com/NeuralNetworkVerification/Marabou. The artifact associated with this tool description is archived on Zenodo [71].
3 Highlighted Features and Applications
In terms of performance, Marabou is on par with state-of-the-art verification tools. In the latest VNN-COMP [11], Marabou won the second place overall, and scored the highest among all CPU-based verifiers. We summarize the main results in Appendix 0.B. In this section, we focus on the usability aspect of Marabou, and highlight some of its recent applications — as well as the features that make them possible. We believe this diverse set of use cases (as well as the relevant scripts in the artifact [71]) serve as valuable examples, which will inspire new ways to apply the solver. A runtime evaluation of Marabou 2.0 against an early version appears in Section 4.
3.0.1 Verifying the Decima job scheduler.
Recently, Graph Neural Networks (GNNs) have been used to schedule jobs over multi-user, distributed-computing clusters, achieving state-of-the-art job completion time [47]. However, concerns remain over whether GNN-based solutions satisfy expected cost-critical properties beyond performance. Marabou has been used to verify a well-known fairness property called strategy-proofness [70] for a high-profile, state-of-the-art GNN-based scheduler called Decima [47]. The verified property states that “a user cannot get their job scheduled earlier by misrepresenting their resource requirement.” While it is challenging to represent a GNN directly in ONNX [25], Marabou’s Python API makes it possible to manually encode Decima and the specification as a set of linear and non-linear constraints. From these constraints, the Network-level Reasoner is able to automatically infer a feed-forward structure with residual connections and then use it for the purpose of abstract interpretation. Notably, Marabou was able to handle the original Decima architecture, proving that the property holds on the vast majority of the examined job profiles but can indeed be violated in some cases.
3.0.2 Formal XAI.
Despite their prevalence, DNNs are considered “black boxes”, uninterpretable to humans. Explainable AI (XAI) aims to understand DNN decisions to enhance trust. Most XAI methods are heuristic-based and lack formal correctness guarantees [54, 45, 55], which can be problematic for critical, regulation-heavy systems. Recent work showed that Marabou can be utilized as a sub-routine in procedures designed for producing formal and provable explanations for DNNs [73, 7, 8, 46, 32]. For instance, it can be used in constructing formal abductive explanations [35, 8], which are subsets of input features that are, by themselves, provably sufficient for determining the DNN’s output. This approach has been successfully applied to large DNNs in the domains of computer vision [73, 8], NLP [46], and DRL robotic navigation [7]. These studies highlight the potential of Marabou in tasks that go beyond formal verification.
3.0.3 Analyzing learning-based robotic systems.
Deep Reinforcement Learning has extensive application in robotic planning and control. Marabou has been applied in these settings to analyze different safety and liveness properties [17, 3, 2, 65]. For example, Amir et al. [2] used Marabou to detect infinite loops in a real-world robotic navigation platform. This was achieved by querying whether there exists a state to which the robot will always return within a finite number of steps $k$, effectively entering an infinite loop. A multi-step property like this can be conveniently encoded in Marabou, by (i) encoding $k$ copies of the control policy; (ii) for each time-step $t$, encoding the system transition as constraints over the current state (input to the policy at $t$), the decided action (output of the policy at $t$), and the next state (input to the policy at $t+1$); and (iii) encoding the “loop” constraint that the initial state ($t_{1}$) is equal to the final state ($t_{k}$). From this set of constraints, the Network-level Reasoner can infer the structure of and perform abstract interpretations over a concatenated network, where the input is the initial state and the output is the final state. Moreover, due to the low input dimension, the split-and-conquer mode in the Multi-thread Manager can be used to perform input-splitting, effectively searching for such loops in independent input regions in parallel. Notably, Marabou can detect loops in the system for agents trained using state-of-the-art RL algorithms, in cases where gradient/optimization-based approaches fail to find any. Loops detected this way have also been observed in the real world [1].
3.0.4 Proof production for the ACAS-Xu benchmarks.
A well-studied set of benchmarks in DNN verification derives from an implementation of the ACAS-Xu airborne system for collision avoidance [38]. Using Marabou, we were able to produce certificates of unsatisfiability for these benchmarks for the first time. Marabou was able to produce certificates for 113 out of the 180 tested benchmarks, with only mild overhead incurred by proof generation and certification. The proof certificates contained over 1.46 million proof-tree leaves, of which more than $99.99\%$ were certified by the native proof checker, while the remaining were certified by a trusted SMT solver. Additional details are provided in [36].
3.0.5 Specifications on neural activation patterns.
Properties of hidden neurons garner increasing interest [77], as they shed light on the internal decision-making process of the neural network. Gopinath et al. [23] observed that for a fixed neural network, certain neuron activation patterns (NAPs) empirically entail a fixed prediction. More recently, Geng et al. [20] formally verified (using Marabou) the aforementioned property, along with a variety of other properties related to NAPs. Specifications related to NAPs can be conveniently encoded in Marabou. For example, specifying that a certain ReLU is activated amounts to setting the lower bound of the variable corresponding to the ReLU input to 0, using the general constraint-encoding methods in the Python/C++ API. Constraints on internal neurons, as with other constraints, can be propagated by the Preprocessor and Network-level Reasoner to tighten bounds.
3.0.6 Robustness against semantically meaningful perturbations.
Considering specifications of perception networks, there is an ongoing effort in the verification community to go beyond adversarial robustness [5, 41, 48, 72, 49]. Marabou has been used to verify robustness against semantically meaningful perturbations that can be analytically defined/abstracted as linear constraints on the neural network inputs (e.g., brightness, uniform haze) [52]. More recently, Marabou has also been successfully applied in a neural symbolic approach, where the correct network behavior is defined with respect to that of another network [74, 72]. For example, Wu et al. [72] considered the specification that an image classifier’s prediction does not change with respect to outputs of an image generative model trained to capture a complex distribution shift (e.g., change in weather condition). A property like this can be conveniently defined in Marabou by loading the classifier and the generator through the Python API and adding the relevant constraints on/between their input and output variables.
3.0.7 Minimal modifications of neural networks.
Recently, multiple papers have proposed techniques to repair a neural network (e.g., by altering selected weights or adding patches) against input points that result in buggy behaviors. Marabou has also been applied in this context to obtain provably minimal modification to the DNN’s weights so that a previously mispredicted input would become correctly predicted. We refer the reader to [53, 22] for additional details.
3.0.8 Verification of quantized neural Networks.
Quantization [37] has become a standard technique for deploying deep neural networks on resource-constrained hardware. The process involves replacing floating point arithmetic with integer arithmetic. Recently, Huang, et.al. [31] have leveraged Marabou to verify quantized neural networks (QNNs). While mainstream platforms like PyTorch do not support saving QNN models in the standard ONNX format, Marabou’s flexible Python API makes it possible to manually encode the quantization operations with constraints such as Round and Clip.
3.0.9 Pruning and slicing neural networks.
Another well-studied approach to reduce the computational cost of deep neural networks is pruning [28, 34], which involves identifying and removing neurons that contribute minimally to the network’s output. Most existing approaches are heuristic-based and lack formal guarantees that the simplified network preserves the behavior of the original. In contrast, it has been shown that Marabou can be used as a sub-procedure in a pruning algorithm [21, 42], which is able to reduce the size of DNNs while provably preserving the output of the neural network up to a given error margin.
3.0.10 Integrating Marabou into complex system verification.
There is a growing demand to embed neural network verification into verification of complex systems that rely on neural networks for perception or decision-making [51]. There are multiple challenges in verifying such systems, and we mention two of them here: (i) currently available provers struggle to overcome the embedding gap, which arises because complex system properties are stated using the abstractions of the problem domain, whereas neural network propeties are stated at a low-level, in terms of normalized real vectors; and (ii) more generally, higher-order theorem provers like HOL, Coq, and Agda can prove more abstract properties of complex systems, but fail at neural network verification; and vice versa: neural network verifiers are not designed to reason about complex hybrid systems. The domain-specific language Vehicle [15, 14] aims at bridging this gap; and includes Marabou as its default verifier. This choice was made thanks to Marabou’s syntax, which is richer and more flexible than available in other, more specialized neural network verifiers.
4 Runtime Evaluation
Figure 3: Runtime performance of Marabou 2.0 and an early version of Marabou on four applications supported by both versions.
We measure the performance improvement in Marabou 2.0 by comparing it against an early Marabou version (git commit 1c1c66), which can handle ReLU and Max constraints and supports symbolic bound propagation [66]. We collected four benchmark sets from the applications described in Section  3: Alternating Loop [2], DeepCert [52], NAP [23, 20], and VeriX [73]. There are 745 instances in total. Details about the benchmarks can be found in Appendix 0.C.1.
Figure 3 compares the runtime of the two Marabou versions on all the benchmarks with a 1 hour CPU timeout. Each configuration was given 1 core and 8GB of memory. Note that Marabou 2.0 was not configured with external solvers in this experiment. We see that Marabou 2.0 is significantly more efficient for a vast majority of the instances. Upon closer examination, an at-least 2×2$\times$2 × speed-up is achieved on 428 instances and an at-least 10×10$\times$10 × speed-up is achieved on 263 instances. Moreover, Marabou 2.0 is also significantly more memory efficient, with a median peak usage of 57MB (versus 604MB with the old version). Solvers’performance on individual benchmarks is reported in Appendix 0.C.2.
5 Conclusion and Next Steps
We have summarized the current state of Marabou, a maturing formal analyzer for neural-network-enabled systems that is under active development. In its current form, Marabou is a versatile and user-friendly toolkit suitable for a wide range of formal analysis tasks. Moving forward, we plan to improve Marabou in several dimensions. Currently, we are actively integrating a CDCL mechanism in the SMT Solver module. Given that many applications involve repeated invocation of the solver on similar queries, we also plan to support incremental solving in the style of pushing and popping constraints, leveraging the newly introduced context-dependent data structures. In addition, adding GPU support (in the Network-level Reasoner) and handling other types of non-linear constraints are also on the development agenda for Marabou.
5.0.1 Acknowledgment
The work of Wu, Zeljić, Tagomori, Huang and Wu was partially supported by the NSF (grant number 2211505), by the BSF (grant number 2020250), a Ford Alliance Project (199909), the Stanford Center for AI Safety, and the Stanford Institute for Human-Centered Artificial Intelligence (HAI). The work of Daggit, Kokke and Komendantskaya was partially supported by the EPSRC grant EP/T026952/1, AISEC: AI Secure and Explainable by Construction. The work of Isac, Refaeli, Amir, Bassan, Lahav and Katz was partially funded by the ISF (grant number 3420/21), by the BSF (grant numbers 2021769 and 2020250), and by the European Union (ERC, VeriDeL, 101112713). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of Zhang was partially supported by the NSFC (grant number 62161146001).
Appendix 0.A Architecture and Core Components
0.A.1 Supported Non-linear Constraints in Marabou 2.0
Type Preproc. NLR SMT (MI)LP CEGAR Proof
ReLU
Max
DNF
Sign
Absolute value
Leaky ReLU
Round
Clip
Sigmoid
Tanh
Softmax
Bi-linear
Appendix 0.B Marabou in VNN-COMP 2023 [11]
Marabou participated in the Fourth International Verification of Neural Networks Competition (VNN-COMP’23) and won the second place overall and scored the highest among all CPU-based verifiers. Table 1 shows the total number of instances solved by the participating solvers. Detailed information can be found in the VNN-COMP 2023 report [11]. The participating version is publicly available in the Marabou repo (commit 1a3ca6). Overall, on the competition benchmarks, there is a non-trivial performance gap between Marabou and GPU-based solvers like $\alpha$-$\beta$-CROWN, especially for larger networks. Supporting GPUs in the Network-level Reasoner is on the development agenda for Marabou.
Table 1: Total number of solved instances in VNN-COMP’23
# Tool Count GPU Support
1 $\alpha$-$\beta$-CROWN 721
2 Marabou 594
3 NeuralSAT 452
4 PyRAT 416
5 nnenum 319
6 NNV 155
7 FastBATLLNN 32
Appendix 0.C Runtime Evaluation
0.C.1 Descriptions of the Benchmarks
The DeepCert benchmarks come from a study [52] of the robustness of image classifiers against a set of analytically defined contextually relevant image perturbations. In particular, we focus on the haze perturbation on an image classifier (model 2A) for the GTSRB dataset. The Alternating Loop benchmarks come from a study of the existence of infinite loops in a robot navigation system [2]. The NAP benchmarks are less straightforward to procure: each benchmark considers an NAP mined on one of the ACAS-Xu networks by Gopinath et al. [23] and checks the robustness of that NAP as defined by Geng et al. [20]. We hope to also evaluate Marabou on the NAP robustness benchmarks described in the latter work, but they are not yet publicly available. Finally, the VeriX benchmarks involve performing a formal explanation extraction procedure (which repeatedly queries Marabou) on a MNIST classifier (the same one used in the original work).
0.C.2 Solver Performance on Individual Benchmarks
Marabou 2.0 Marabou (1c1c66)
Family Benchmark Result Time Mem. Result Time Mem.
[…]
Alt. Loops REI.id114.ep84297 UNSAT 55.7 10.6 UNSAT 1036.7 50.9
Alt. Loops REI.id114.ep94159 SAT 0.6 6.7 SAT 22.3 42.6
Alt. Loops REI.id114.ep94581 SAT 0.6 6.7 SAT 35.9 46.2
Alt. Loops REI.id114.ep97993 SAT 0.5 6.7 SAT 139.3 48.4
Alt. Loops REI.id128.ep68334 UNSAT 111.0 9.6 UNSAT 1749.2 56.0
Alt. Loops REI.id128.ep85878 UNSAT 171.1 10.2 UNSAT 3478.2 60.4
Alt. Loops REI.id128.ep87384 SAT 0.7 6.3 SAT 308.5 57.1
Alt. Loops REI.id128.ep91042 UNSAT 94.9 9.2 UNSAT 1267.4 53.9
Alt. Loops REI.id128.ep97983 UNSAT 127.2 10.0 UNSAT 2404.6 56.7
Alt. Loops REI.id158.ep53113 SAT 70.9 9.7 SAT 16.9 34.8
Alt. Loops REI.id158.ep53499 SAT 14.6 8.6 SAT 34.1 36.8
Alt. Loops REI.id158.ep53554 SAT 10.4 8.9 SAT 17.7 39.0
Alt. Loops REI.id158.ep59378 UNSAT 32.0 9.8 UNSAT 80.8 37.6
Alt. Loops REI.id158.ep63427 UNSAT 23.7 9.4 UNSAT 68.9 38.7
Alt. Loops REI.id165.ep82822 SAT 0.7 6.8 SAT 13.2 31.6
Alt. Loops REI.id165.ep94134 SAT 0.7 6.7 SAT 14.4 33.9
Alt. Loops REI.id165.ep94522 SAT 0.5 6.5 SAT 0.2 11.2
Alt. Loops REI.id165.ep95976 SAT 0.5 6.8 SAT 30.7 34.6
Alt. Loops REI.id165.ep96974 SAT 24.0 8.3 SAT 16.8 32.5
Alt. Loops REI.id174.ep66258 SAT 10.5 9.6 SAT 0.1 5.8
Alt. Loops REI.id174.ep81057 SAT 0.6 6.5 SAT 0.3 19.6
Alt. Loops REI.id174.ep91107 SAT 18.0 9.5 SAT 49.7 43.4
Alt. Loops REI.id174.ep97933 SAT 0.7 6.7 SAT 10.6 41.8
Alt. Loops REI.id174.ep98455 SAT 25.5 10.3 SAT 0.2 13.0
Alt. Loops REI.id176.ep92206 UNSAT 507.3 10.8 UNK 3601.0 66.0
Alt. Loops REI.id176.ep95516 UNSAT 438.3 10.4 UNK 3601.0 61.5
Alt. Loops REI.id176.ep96138 UNSAT 364.7 10.1 UNK 3601.0 63.1
Alt. Loops REI.id176.ep98430 UNSAT 816.4 10.8 UNK 3601.0 64.6
Alt. Loops REI.id176.ep99427 UNSAT 410.6 10.3 UNK 3601.0 64.2
Alt. Loops REI.id180.ep58329 SAT 0.5 6.2 SAT 0.3 15.8
Alt. Loops REI.id180.ep58461 SAT 0.5 6.4 SAT 0.2 14.3
Alt. Loops REI.id180.ep61293 SAT 0.6 6.4 SAT 0.6 28.1
Alt. Loops REI.id180.ep69795 SAT 0.5 6.4 SAT 0.2 9.6
Alt. Loops REI.id180.ep77624 SAT 0.4 6.6 SAT 0.1 9.3
Alt. Loops REI.id196.ep41633 SAT 0.8 6.6 SAT 0.2 12.0
Alt. Loops REI.id196.ep43251 SAT 0.3 6.4 SAT 0.1 9.3
Alt. Loops REI.id196.ep73462 SAT 0.7 6.7 SAT 0.2 12.8
Alt. Loops REI.id196.ep86142 SAT 0.3 6.4 SAT 0.1 8.3
Alt. Loops REI.id196.ep99092 SAT 0.4 6.2 SAT 24.0 34.0
Alt. Loops REI.id201.ep39282 SAT 21.5 9.6 SAT 4.0 39.8
Alt. Loops REI.id201.ep48186 SAT 0.6 6.2 SAT 51.4 43.1
Alt. Loops REI.id201.ep48219 SAT 0.6 6.4 SAT 0.2 10.6
Alt. Loops REI.id201.ep56910 SAT 0.4 6.4 SAT 28.7 34.9
Alt. Loops REI.id201.ep88825 SAT 0.2 6.0 SAT 49.6 45.4
Alt. Loops REI.id215.ep96277 SAT 0.5 6.6 SAT 1.7 41.1
Alt. Loops REI.id215.ep97333 SAT 0.3 6.4 SAT 4.4 40.8
Alt. Loops REI.id215.ep97355 SAT 0.6 6.4 SAT 0.6 32.9
Alt. Loops REI.id215.ep97421 SAT 0.5 6.9 SAT 1763.0 60.1
Alt. Loops REI.id215.ep98767 SAT 0.5 6.5 SAT 2.1 35.0
Alt. Loops REI.id234.ep70556 SAT 0.6 6.5 SAT 0.1 10.8
Alt. Loops REI.id234.ep71035 SAT 0.3 6.2 SAT 0.3 24.4
Alt. Loops REI.id234.ep71965 SAT 0.6 7.1 SAT 0.2 11.8
Alt. Loops REI.id234.ep83703 SAT 0.7 6.7 SAT 0.3 22.4
Alt. Loops REI.id234.ep83780 SAT 0.4 6.4 SAT 1.0 33.8
Alt. Loops REI.id239.ep71386 UNSAT 209.9 10.2 UNSAT 1691.7 55.5
Alt. Loops REI.id239.ep73100 UNSAT 210.3 10.4 UNSAT 1110.3 60.0
Alt. Loops REI.id239.ep73504 SAT 6.7 9.1 SAT 3.4 51.2
Alt. Loops REI.id239.ep79112 UNSAT 193.2 9.9 UNSAT 1643.7 56.7
Alt. Loops REI.id239.ep90118 SAT 35.8 9.3 SAT 0.4 18.0
Alt. Loops REI.id240.ep85050 UNSAT 37.2 9.7 UNSAT 308.5 47.0
Alt. Loops REI.id240.ep89317 UNSAT 26.5 9.4 UNSAT 152.7 42.7
Alt. Loops REI.id240.ep91714 UNSAT 28.2 9.9 UNSAT 177.1 42.5
Alt. Loops REI.id240.ep98115 UNSAT 40.2 10.2 UNSAT 282.4 47.6
Alt. Loops REI.id240.ep98891 UNSAT 28.1 9.6 UNSAT 76.3 45.4
Alt. Loops REI.id267.ep91426 UNSAT 0.7 6.2 UNSAT 0.7 14.5
Alt. Loops REI.id267.ep95713 UNSAT 1.7 6.4 UNSAT 2.5 16.7
Alt. Loops REI.id267.ep96300 UNSAT 1.7 6.9 UNSAT 1.0 14.8
Alt. Loops REI.id267.ep96333 UNSAT 1.1 6.2 UNSAT 0.5 12.4
Alt. Loops REI.id267.ep96426 UNSAT 1.3 6.0 UNSAT 1.5 15.5
Alt. Loops REI.id286.ep89988 UNSAT 32.9 8.5 UNSAT 41.8 35.9
Alt. Loops REI.id286.ep89999 UNSAT 32.4 8.5 UNSAT 41.7 35.8
Alt. Loops REI.id286.ep90032 UNSAT 39.1 9.0 UNSAT 59.7 37.9
Alt. Loops REI.id286.ep94925 UNSAT 35.1 9.2 UNSAT 56.8 38.1
Alt. Loops REI.id286.ep95719 UNSAT 50.8 9.0 UNSAT 63.4 38.2
Alt. Loops REI.id296.ep74854 UNSAT 140.7 9.7 UNSAT 697.8 51.9
Alt. Loops REI.id296.ep74969 UNSAT 119.7 9.6 UNSAT 559.8 55.9
Alt. Loops REI.id296.ep76414 UNSAT 127.1 10.1 UNSAT 500.1 51.7
Alt. Loops REI.id296.ep77867 UNSAT 100.3 9.4 UNSAT 131.1 48.4
Alt. Loops REI.id296.ep79125 UNSAT 133.1 9.9 UNSAT 274.0 46.4
Alt. Loops REI.id298.ep75768 UNSAT 240.6 10.9 UNK 3601.0 66.3
Alt. Loops REI.id298.ep77187 SAT 0.8 6.8 SAT 0.4 36.5
Alt. Loops REI.id298.ep79143 UNSAT 167.6 10.9 UNK 3601.0 70.0
Alt. Loops REI.id298.ep85589 UNSAT 175.5 10.8 UNK 3601.0 65.9
Alt. Loops REI.id298.ep97154 SAT 0.5 6.8 SAT 0.1 8.0
Alt. Loops REI.id299.ep69103 UNSAT 57.5 11.2 UNSAT 479.8 49.8
Alt. Loops REI.id299.ep69400 UNSAT 49.6 10.7 UNSAT 361.7 47.0
Alt. Loops REI.id299.ep81175 UNSAT 40.3 9.7 UNSAT 284.3 45.7
Alt. Loops REI.id299.ep92705 UNSAT 111.3 10.7 UNSAT 739.8 50.7
Alt. Loops REI.id299.ep99426 UNSAT 69.2 9.8 UNSAT 569.1 49.5
Alt. Loops REI.id303.ep68520 UNSAT 243.1 11.6 UNK 3601.0 64.6
Alt. Loops REI.id303.ep69708 SAT 1.6 8.6 SAT 987.6 61.3
Alt. Loops REI.id303.ep94994 SAT 0.6 6.6 SAT 72.8 58.6
Alt. Loops REI.id303.ep95038 SAT 0.6 6.4 SAT 4.4 52.3
Alt. Loops REI.id303.ep95137 SAT 0.4 6.4 SAT 1399.1 72.3
Alt. Loops REI.id308.ep94787 UNSAT 19.9 8.7 UNSAT 95.0 46.1
Alt. Loops REI.id308.ep97397 SAT 0.5 6.7 SAT 0.2 14.4
Alt. Loops REI.id308.ep98363 SAT 0.7 6.6 SAT 125.7 45.1
Alt. Loops REI.id308.ep98830 UNSAT 52.5 10.1 UNSAT 207.0 46.9
Alt. Loops REI.id308.ep99303 SAT 0.8 6.9 SAT 6.1 35.9
Alt. Loops REI.id318.ep75036 SAT 0.6 6.3 SAT 14.4 64.8
Alt. Loops REI.id318.ep85341 SAT 96.4 10.7 SAT 164.6 57.5
Alt. Loops REI.id318.ep89536 SAT 0.7 6.5 SAT 27.3 55.8
Alt. Loops REI.id318.ep94480 SAT 0.5 6.5 SAT 0.9 37.7
Alt. Loops REI.id318.ep96308 SAT 0.6 6.7 SAT 0.4 21.7
Alt. Loops REI.id319.ep62661 SAT 0.4 6.5 SAT 71.0 54.8
Alt. Loops REI.id319.ep84434 SAT 0.6 6.9 SAT 2437.7 67.5
Alt. Loops REI.id319.ep84478 SAT 0.3 6.4 SAT 0.3 14.3
Alt. Loops REI.id319.ep86630 SAT 0.5 6.2 SAT 573.1 62.2
Alt. Loops REI.id319.ep89649 SAT 0.3 6.6 SAT 0.3 12.9
Alt. Loops REI.id321.ep55647 UNSAT 9.7 8.5 UNSAT 19.9 33.0
Alt. Loops REI.id321.ep81083 UNSAT 18.2 8.5 UNSAT 404.1 53.6
Alt. Loops REI.id321.ep86834 UNSAT 28.6 9.3 UNSAT 712.3 62.0
Alt. Loops REI.id321.ep92994 UNSAT 25.0 9.5 UNSAT 389.5 54.3
Alt. Loops REI.id321.ep93439 UNSAT 40.2 8.9 UNSAT 686.8 58.2
Alt. Loops REI.id343.ep55239 UNSAT 62.6 9.2 UNSAT 149.7 47.1
Alt. Loops REI.id343.ep63040 UNSAT 83.4 9.1 UNSAT 168.2 45.4
Alt. Loops REI.id343.ep63051 UNSAT 83.2 9.1 UNSAT 166.6 45.5
Alt. Loops REI.id343.ep73171 UNSAT 156.1 9.7 UNSAT 1022.7 54.2
Alt. Loops REI.id343.ep96195 UNSAT 235.0 9.3 UNSAT 3113.7 57.7
Alt. Loops REI.id35.ep61486 SAT 0.4 6.4 SAT 22.1 39.1
Alt. Loops REI.id35.ep61788 SAT 0.5 6.7 SAT 0.2 10.9
Alt. Loops REI.id35.ep81067 SAT 0.3 6.3 SAT 0.2 12.3
Alt. Loops REI.id35.ep85842 SAT 0.2 6.1 SAT 0.2 10.4
Alt. Loops REI.id35.ep94624 SAT 0.4 6.4 SAT 0.1 7.1
Alt. Loops REI.id352.ep49376 UNSAT 7.6 8.2 UNSAT 58.8 39.3
Alt. Loops REI.id352.ep51510 UNSAT 8.0 9.0 UNSAT 169.9 43.4
Alt. Loops REI.id352.ep68171 UNSAT 12.1 9.2 UNSAT 87.1 39.8
Alt. Loops REI.id352.ep69872 UNSAT 9.5 8.8 UNSAT 68.1 38.3
Alt. Loops REI.id352.ep70312 UNSAT 8.1 8.2 UNSAT 24.0 37.8
Alt. Loops REI.id379.ep71969 SAT 1.4 7.9 SAT 13.7 43.2
Alt. Loops REI.id379.ep80359 SAT 0.7 6.8 SAT 0.2 11.7
Alt. Loops REI.id379.ep87708 UNSAT 212.5 10.6 UNSAT 2425.8 57.3
Alt. Loops REI.id379.ep96909 SAT 32.1 10.7 SAT 180.9 69.4
Alt. Loops REI.id379.ep97133 SAT 1.2 7.2 SAT 0.7 36.4
Alt. Loops REI.id381.ep85470 SAT 1.9 9.2 UNK 3601.0 64.0
Alt. Loops REI.id381.ep85558 SAT 1.7 8.8 SAT 1140.9 62.2
Alt. Loops REI.id381.ep92541 UNSAT 250.5 11.2 UNK 3600.9 59.8
Alt. Loops REI.id381.ep92994 SAT 1.3 7.3 UNK 3601.0 64.2
Alt. Loops REI.id381.ep93562 UNSAT 281.6 10.7 UNK 3601.0 64.3
Alt. Loops REI.id393.ep88789 SAT 25.5 11.0 SAT 0.3 17.0
Alt. Loops REI.id393.ep90550 SAT 0.5 6.4 SAT 64.9 56.5
Alt. Loops REI.id393.ep97111 SAT 31.6 10.0 SAT 0.1 10.0
Alt. Loops REI.id393.ep97199 SAT 43.0 10.6 SAT 0.2 15.2
Alt. Loops REI.id393.ep97221 SAT 37.3 11.0 SAT 0.2 14.3
Alt. Loops REI.id399.ep62611 UNSAT 42.9 9.2 UNSAT 490.9 56.3
Alt. Loops REI.id399.ep69315 UNSAT 43.9 9.8 UNSAT 3342.7 68.6
Alt. Loops REI.id399.ep85537 UNSAT 91.9 9.5 UNSAT 2220.3 60.8
Alt. Loops REI.id399.ep92145 UNSAT 47.1 8.8 UNSAT 1475.9 64.6
Alt. Loops REI.id399.ep96449 SAT 1.5 7.1 SAT 140.4 64.8
Alt. Loops REI.id418.ep91524 SAT 0.4 6.2 SAT 5.3 53.1
Alt. Loops REI.id418.ep92561 SAT 0.6 6.6 SAT 19.8 51.7
Alt. Loops REI.id418.ep92671 SAT 0.5 6.4 SAT 30.0 49.7
Alt. Loops REI.id418.ep99126 SAT 0.9 6.4 SAT 134.2 51.4
Alt. Loops REI.id418.ep99947 SAT 0.6 6.5 SAT 13.4 51.0
Alt. Loops REI.id425.ep69300 SAT 0.6 6.7 SAT 0.8 37.0
Alt. Loops REI.id425.ep83340 SAT 0.6 6.9 SAT 0.3 14.0
Alt. Loops REI.id425.ep93745 UNSAT 113.7 10.0 UNSAT 1299.1 51.4
Alt. Loops REI.id425.ep96981 SAT 0.5 6.6 SAT 0.1 8.0
Alt. Loops REI.id425.ep98177 SAT 0.5 6.6 SAT 0.1 11.8
Alt. Loops REI.id444.ep77347 UNSAT 90.7 10.7 UNSAT 501.8 50.7
Alt. Loops REI.id444.ep78487 UNSAT 120.5 10.2 UNSAT 1094.5 51.6
Alt. Loops REI.id444.ep87148 UNSAT 78.3 10.0 UNSAT 842.5 51.1
Alt. Loops REI.id444.ep92635 UNSAT 105.6 10.1 UNK 3601.0 48.8
Alt. Loops REI.id444.ep96160 UNSAT 89.3 9.7 UNSAT 1262.5 51.4
Alt. Loops REI.id457.ep46821 SAT 11.0 9.3 SAT 0.1 8.8
Alt. Loops REI.id457.ep69348 SAT 0.5 6.9 SAT 0.2 10.6
Alt. Loops REI.id457.ep78403 SAT 0.3 6.4 SAT 0.1 7.3
Alt. Loops REI.id457.ep79411 SAT 0.5 6.3 SAT 0.2 12.7
Alt. Loops REI.id457.ep81404 SAT 0.6 6.5 SAT 0.1 8.5
Alt. Loops REI.id47.ep79665 SAT 0.8 6.6 SAT 0.3 19.2
Alt. Loops REI.id47.ep92041 SAT 0.3 6.3 SAT 0.1 10.0
Alt. Loops REI.id47.ep96122 SAT 0.7 6.6 SAT 0.4 21.0
Alt. Loops REI.id47.ep99600 SAT 0.6 6.4 SAT 0.3 18.8
Alt. Loops REI.id47.ep99688 SAT 0.5 7.1 SAT 0.3 18.2
Alt. Loops REI.id491.ep54571 SAT 0.3 6.6 SAT 0.6 25.5
Alt. Loops REI.id491.ep67504 SAT 0.6 6.7 SAT 1.3 33.2
Alt. Loops REI.id491.ep67924 SAT 0.4 6.6 SAT 0.2 12.8
Alt. Loops REI.id491.ep70134 SAT 0.5 6.5 SAT 0.4 23.2
Alt. Loops REI.id491.ep96326 SAT 0.4 6.1 SAT 5.6 39.0
Alt. Loops REI.id502.ep71027 SAT 0.4 6.5 SAT 56.7 42.9
Alt. Loops REI.id502.ep71445 SAT 0.4 6.5 SAT 15.9 46.3
Alt. Loops REI.id502.ep84337 SAT 0.6 7.0 SAT 2.1 30.2
Alt. Loops REI.id502.ep86449 SAT 0.8 6.9 SAT 21.3 42.0
Alt. Loops REI.id502.ep90122 SAT 0.4 6.4 SAT 266.6 45.8
Alt. Loops REI.id512.ep70217 SAT 0.6 6.6 SAT 1.3 41.2
Alt. Loops REI.id512.ep70283 SAT 0.6 6.8 SAT 1.0 33.3
Alt. Loops REI.id512.ep75804 SAT 0.5 6.4 SAT 986.8 50.4
Alt. Loops REI.id512.ep97330 SAT 0.6 6.8 SAT 0.1 11.2
Alt. Loops REI.id512.ep97385 SAT 0.4 6.3 SAT 567.2 47.5
Alt. Loops REI.id518.ep68431 UNSAT 97.2 9.9 UNSAT 523.8 46.8
Alt. Loops REI.id518.ep70223 UNSAT 85.6 9.2 UNSAT 1358.8 49.7
Alt. Loops REI.id518.ep95639 UNSAT 125.9 10.2 UNSAT 132.8 39.7
Alt. Loops REI.id518.ep95900 UNSAT 123.8 9.9 UNSAT 602.4 45.1
Alt. Loops REI.id518.ep96115 UNSAT 89.9 9.9 UNSAT 561.7 45.2
Alt. Loops REI.id528.ep55711 SAT 1.4 7.4 SAT 1446.3 51.4
Alt. Loops REI.id528.ep89177 SAT 0.9 6.5 SAT 53.8 53.1
Alt. Loops REI.id528.ep89331 SAT 0.8 6.9 SAT 158.9 48.1
Alt. Loops REI.id528.ep90341 SAT 1.0 6.6 SAT 36.3 39.8
Alt. Loops REI.id528.ep95627 SAT 0.5 6.3 SAT 0.1 8.3
Alt. Loops REI.id530.ep82130 SAT 0.5 6.4 SAT 0.2 13.2
Alt. Loops REI.id530.ep82185 SAT 0.5 6.7 SAT 0.3 12.9
Alt. Loops REI.id530.ep82207 SAT 0.5 6.4 SAT 0.2 13.1
Alt. Loops REI.id530.ep91636 SAT 0.5 6.8 SAT 0.4 31.3
Alt. Loops REI.id530.ep94463 SAT 0.6 6.7 SAT 0.2 12.8
Alt. Loops REI.id535.ep76894 UNSAT 22.1 9.7 UNSAT 54.5 43.6
Alt. Loops REI.id535.ep76916 UNSAT 58.5 9.3 UNSAT 51.4 39.6
Alt. Loops REI.id535.ep87193 SAT 0.8 6.8 SAT 0.2 13.9
Alt. Loops REI.id535.ep87226 SAT 0.6 6.9 SAT 0.5 22.8
Alt. Loops REI.id535.ep91323 UNSAT 16.7 9.4 UNSAT 176.3 44.0
Alt. Loops REI.id540.ep57595 UNSAT 1415.4 10.5 UNK 3601.0 65.2
Alt. Loops REI.id540.ep57610 UNSAT 1984.2 10.8 UNK 3601.0 66.9
Alt. Loops REI.id540.ep57775 UNSAT 1625.2 10.5 UNK 3601.0 68.2
Alt. Loops REI.id540.ep85904 SAT 0.8 6.8 SAT 762.1 68.3
Alt. Loops REI.id540.ep94690 SAT 0.5 6.2 SAT 0.1 12.2
Alt. Loops REI.id549.ep85789 SAT 0.5 6.6 SAT 0.2 15.2
Alt. Loops REI.id549.ep87887 SAT 506.8 10.8 SAT 1509.6 64.1
Alt. Loops REI.id549.ep93366 SAT 0.4 6.4 SAT 28.6 51.1
Alt. Loops REI.id549.ep99098 SAT 1.0 6.8 SAT 0.3 16.4
Alt. Loops REI.id549.ep99331 SAT 0.6 6.5 SAT 0.1 8.2
Alt. Loops REI.id67.ep82981 UNSAT 27.2 10.2 UNSAT 48.7 33.4
Alt. Loops REI.id67.ep84585 UNSAT 53.2 10.6 UNSAT 98.8 39.6
Alt. Loops REI.id67.ep87611 UNSAT 50.7 11.0 UNSAT 29.7 34.7
Alt. Loops REI.id67.ep95115 SAT 0.6 6.6 SAT 111.2 41.5
Alt. Loops REI.id67.ep99214 SAT 0.6 6.8 SAT 0.8 35.5
Alt. Loops REI.id68.ep95777 SAT 0.5 6.8 SAT 0.4 20.0
Alt. Loops REI.id68.ep95821 SAT 0.6 6.5 SAT 0.2 10.8
Alt. Loops REI.id68.ep95854 SAT 0.6 6.8 SAT 0.3 16.7
Alt. Loops REI.id68.ep95942 SAT 0.5 6.4 SAT 0.2 11.4
Alt. Loops REI.id68.ep98804 SAT 0.5 6.6 SAT 0.3 23.0
Alt. Loops REI.id73.ep74414 SAT 328.7 11.0 SAT 81.0 52.7
Alt. Loops REI.id73.ep74469 SAT 308.8 11.4 SAT 23.7 50.4
Alt. Loops REI.id73.ep82370 UNSAT 221.5 11.0 UNSAT 1166.3 58.1
Alt. Loops REI.id73.ep84023 UNSAT 137.7 10.5 UNSAT 946.3 54.9
Alt. Loops REI.id73.ep91504 SAT 0.4 6.2 SAT 153.7 53.7
Alt. Loops REI.id76.ep90543 UNSAT 216.7 10.3 UNSAT 2095.2 62.3
Alt. Loops REI.id76.ep91480 UNSAT 289.6 10.8 UNSAT 1264.9 62.4
Alt. Loops REI.id76.ep92143 UNSAT 208.5 10.6 UNSAT 411.0 46.4
Alt. Loops REI.id76.ep98692 UNSAT 310.5 11.0 UNSAT 1011.0 54.6
Alt. Loops REI.id76.ep99198 UNSAT 315.0 10.9 UNSAT 1862.4 56.5
Alt. Loops REI.id77.ep49509 UNSAT 25.0 8.8 UNSAT 69.4 37.8
Alt. Loops REI.id77.ep75183 UNSAT 29.6 8.9 UNSAT 102.2 38.4
Alt. Loops REI.id77.ep91446 UNSAT 52.7 9.6 UNSAT 165.0 41.0
Alt. Loops REI.id77.ep99119 UNSAT 79.7 9.7 UNSAT 147.6 38.8
Alt. Loops REI.id77.ep99229 UNSAT 73.5 8.9 UNSAT 144.3 40.5
Alt. Loops REI.id81.ep89983 UNSAT 176.4 10.7 UNK 3601.0 59.4
Alt. Loops REI.id81.ep91408 UNSAT 146.9 10.7 UNK 3601.0 60.9
Alt. Loops REI.id81.ep91571 UNSAT 177.1 10.4 UNSAT 2485.5 56.8
Alt. Loops REI.id81.ep97775 UNSAT 204.1 11.0 UNSAT 3578.0 61.4
Alt. Loops REI.id81.ep97833 UNSAT 186.6 10.6 UNK 3601.0 57.0
Alt. Loops REI.id90.ep84292 UNSAT 318.8 10.2 UNK 3601.0 60.4
Alt. Loops REI.id90.ep94747 UNSAT 285.4 10.6 UNK 3601.0 59.0
Alt. Loops REI.id90.ep97680 UNSAT 174.2 10.3 UNK 3601.0 54.8
Alt. Loops REI.id90.ep97955 UNSAT 167.0 10.2 UNK 3601.0 59.9
Alt. Loops REI.id90.ep99033 UNSAT 202.8 10.3 UNK 3601.0 58.0
Alt. Loops REI.id91.ep54663 UNSAT 46.6 8.7 UNSAT 1790.1 53.0
Alt. Loops REI.id91.ep58150 UNSAT 51.6 10.6 UNK 3601.0 55.4
Alt. Loops REI.id91.ep58172 UNSAT 38.4 9.2 UNSAT 2933.8 56.0
Alt. Loops REI.id91.ep75275 UNSAT 104.4 10.7 UNSAT 649.1 48.4
Alt. Loops REI.id91.ep95105 UNSAT 125.5 10.0 UNSAT 1257.0 56.5
DeepCert ind0.target6.eps0.2 UNSAT 8.7 399.7 UNSAT 268.9 603.7
DeepCert ind0.target6.eps0.4 UNSAT 8.5 412.4 UNSAT 830.0 2053.0
DeepCert ind0.target6.eps0.6 UNSAT 24.8 569.3 UNSAT 770.8 1252.2
DeepCert ind14.target6.eps0.2 UNSAT 8.5 402.8 UNSAT 457.9 1081.1
DeepCert ind14.target6.eps0.4 SAT 26.9 561.2 SAT 617.3 1421.2
DeepCert ind14.target6.eps0.6 SAT 24.3 569.9 SAT 428.7 1086.1
DeepCert ind15.target6.eps0.2 UNSAT 8.5 401.4 UNSAT 398.0 602.4
DeepCert ind15.target6.eps0.4 UNSAT 39.4 560.2 UNSAT 476.5 951.0
DeepCert ind15.target6.eps0.6 SAT 24.4 568.7 SAT 177.8 753.4
DeepCert ind19.target6.eps0.2 UNSAT 8.5 402.8 UNSAT 442.5 938.9
DeepCert ind19.target6.eps0.4 UNSAT 33.0 560.7 UNSAT 532.1 1419.9
DeepCert ind19.target6.eps0.6 SAT 25.5 569.5 SAT 419.0 754.9
DeepCert ind20.target6.eps0.2 UNSAT 8.6 403.3 UNSAT 352.1 603.2
DeepCert ind20.target6.eps0.4 UNSAT 25.8 560.9 UNSAT 1568.0 2244.6
DeepCert ind20.target6.eps0.6 SAT 24.3 570.9 SAT 436.5 1403.0
DeepCert ind21.target6.eps0.2 UNSAT 8.7 402.4 UNSAT 366.8 601.6
DeepCert ind21.target6.eps0.4 UNSAT 32.9 560.4 UNSAT 1054.4 1889.6
DeepCert ind21.target6.eps0.6 SAT 27.9 568.9 SAT 874.6 2369.0
DeepCert ind22.target6.eps0.2 UNSAT 8.6 403.0 UNSAT 261.3 602.2
DeepCert ind22.target6.eps0.4 UNSAT 26.5 561.2 UNSAT 698.6 943.7
DeepCert ind22.target6.eps0.6 SAT 27.3 569.6 SAT 376.1 3149.2
DeepCert ind23.target6.eps0.2 UNSAT 8.8 401.9 UNSAT 615.3 1075.1
DeepCert ind23.target6.eps0.4 UNSAT 33.7 560.4 UNSAT 727.4 1255.8
DeepCert ind23.target6.eps0.6 SAT 25.8 568.0 SAT 162.7 910.2
DeepCert ind24.target6.eps0.2 UNSAT 10.0 401.6 UNSAT 829.5 1259.0
DeepCert ind24.target6.eps0.4 UNSAT 35.1 559.1 UNSAT 1477.3 1929.0
DeepCert ind24.target6.eps0.6 SAT 26.8 567.9 SAT 272.1 933.5
DeepCert ind25.target6.eps0.2 UNSAT 8.5 403.3 UNSAT 933.3 1572.4
DeepCert ind25.target6.eps0.4 SAT 24.6 560.6 SAT 293.6 925.2
DeepCert ind25.target6.eps0.6 SAT 23.7 569.4 SAT 357.4 1251.2
DeepCert ind26.target6.eps0.2 UNSAT 8.6 404.0 UNSAT 515.5 781.7
DeepCert ind26.target6.eps0.4 UNSAT 25.1 561.0 UNSAT 1136.0 1887.4
DeepCert ind26.target6.eps0.6 SAT 24.0 568.7 SAT 1453.1 1568.0
DeepCert ind27.target6.eps0.2 UNSAT 8.6 402.4 UNSAT 500.2 781.4
DeepCert ind27.target6.eps0.4 SAT 24.2 561.1 SAT 628.1 1271.7
DeepCert ind27.target6.eps0.6 SAT 24.1 568.9 SAT 905.0 2196.7
DeepCert ind28.target6.eps0.2 UNSAT 8.5 402.4 UNSAT 714.5 945.7
DeepCert ind28.target6.eps0.4 UNSAT 30.7 561.5 UNSAT 433.7 939.1
DeepCert ind28.target6.eps0.6 SAT 24.8 569.2 SAT 437.0 1892.1
DeepCert ind29.target6.eps0.2 UNSAT 8.4 402.1 UNSAT 597.4 782.6
DeepCert ind29.target6.eps0.4 SAT 23.6 561.2 SAT 687.2 1112.0
DeepCert ind29.target6.eps0.6 SAT 25.2 569.3 SAT 461.1 1248.9
DeepCert ind31.target6.eps0.2 UNSAT 26.4 540.6 UNSAT 549.4 949.7
DeepCert ind31.target6.eps0.4 SAT 23.6 554.4 SAT 309.5 1089.2
DeepCert ind31.target6.eps0.6 SAT 24.1 564.2 SAT 366.0 1268.6
DeepCert ind33.target6.eps0.2 UNSAT 20.9 539.6 UNSAT 804.9 1414.6
DeepCert ind33.target6.eps0.4 SAT 23.1 553.2 SAT 211.3 1409.0
DeepCert ind33.target6.eps0.6 SAT 23.8 564.2 SAT 357.8 757.1
DeepCert ind39.target6.eps0.2 UNSAT 8.3 395.9 UNSAT 300.2 603.1
DeepCert ind39.target6.eps0.4 UNSAT 34.5 553.4 UNSAT 1185.9 3033.1
DeepCert ind39.target6.eps0.6 UNSAT 32.7 562.7 UNSAT 1742.9 2882.3
DeepCert ind42.target6.eps0.2 UNSAT 8.4 396.2 UNSAT 603.8 940.5
DeepCert ind42.target6.eps0.4 UNSAT 37.4 552.2 UNSAT 1143.6 3472.4
DeepCert ind42.target6.eps0.6 UNSAT 56.1 561.8 UNSAT 1596.8 2047.7
DeepCert ind43.target6.eps0.2 UNSAT 8.4 398.1 UNSAT 340.0 602.7
DeepCert ind43.target6.eps0.4 UNSAT 29.6 555.1 UNSAT 1673.5 2244.1
DeepCert ind43.target6.eps0.6 UNSAT 31.8 564.1 UNSAT 1622.2 3670.7
DeepCert ind45.target6.eps0.2 UNSAT 8.4 397.3 UNSAT 428.6 781.9
DeepCert ind45.target6.eps0.4 UNSAT 35.5 552.3 UNSAT 1208.6 2048.1
DeepCert ind45.target6.eps0.6 UNSAT 25.6 562.3 UNSAT 2229.4 3658.6
DeepCert ind46.target6.eps0.2 UNSAT 8.1 394.3 UNSAT 1287.0 1890.3
DeepCert ind46.target6.eps0.4 UNSAT 24.6 548.8 UNSAT 1902.9 2328.9
DeepCert ind46.target6.eps0.6 UNSAT 38.4 559.0 UNSAT 951.6 2042.0
DeepCert ind47.target6.eps0.2 UNSAT 8.1 393.1 UNSAT 631.0 941.5
DeepCert ind47.target6.eps0.4 UNSAT 24.4 547.9 UNSAT 736.4 1258.9
DeepCert ind47.target6.eps0.6 UNSAT 30.7 557.4 UNSAT 1700.8 1771.6
DeepCert ind51.target6.eps0.2 UNSAT 8.2 394.8 UNSAT 365.8 1102.2
DeepCert ind51.target6.eps0.4 UNSAT 24.1 551.1 UNSAT 1807.4 2528.8
DeepCert ind51.target6.eps0.6 UNSAT 40.6 559.9 UNSAT 2099.4 1973.8
DeepCert ind53.target6.eps0.2 UNSAT 8.0 392.8 UNSAT 376.4 783.7
DeepCert ind53.target6.eps0.4 UNSAT 24.5 547.4 UNSAT 1429.3 2523.9
DeepCert ind53.target6.eps0.6 UNSAT 34.4 556.1 UNSAT 1480.0 1731.3
DeepCert ind55.target6.eps0.2 UNSAT 7.9 389.9 UNSAT 235.3 600.9
DeepCert ind55.target6.eps0.4 UNSAT 24.4 544.1 UNSAT 1388.7 1972.8
DeepCert ind55.target6.eps0.6 UNSAT 30.8 553.0 UNSAT 875.4 2053.1
DeepCert ind56.target6.eps0.2 UNSAT 8.0 391.0 UNSAT 579.3 1101.2
DeepCert ind56.target6.eps0.4 UNSAT 23.5 544.3 UNSAT 809.9 1137.7
DeepCert ind56.target6.eps0.6 UNSAT 38.8 553.4 UNSAT 1320.3 1728.9
DeepCert ind6.target6.eps0.2 UNSAT 8.7 400.1 UNSAT 417.5 1098.1
DeepCert ind6.target6.eps0.4 UNSAT 23.9 555.5 UNSAT 701.0 1576.8
DeepCert ind6.target6.eps0.6 SAT 26.1 564.1 SAT 734.7 1255.1
DeepCert ind60.target6.eps0.2 UNSAT 6.2 347.4 UNSAT 543.8 604.8
DeepCert ind60.target6.eps0.4 UNSAT 6.0 358.8 UNSAT 302.9 599.8
DeepCert ind60.target6.eps0.6 UNSAT 24.4 492.7 UNSAT 3366.6 2526.6
DeepCert ind66.target6.eps0.2 UNSAT 5.1 321.8 UNSAT 465.4 601.4
DeepCert ind66.target6.eps0.4 SAT 14.5 443.9 SAT 1079.6 1255.9
DeepCert ind66.target6.eps0.6 SAT 14.8 453.3 SAT 1648.3 2406.2
DeepCert ind67.target6.eps0.2 UNSAT 5.2 323.0 UNSAT 503.2 602.8
DeepCert ind67.target6.eps0.4 UNSAT 18.6 447.2 UNSAT 994.9 1110.3
DeepCert ind67.target6.eps0.6 SAT 16.0 456.6 SAT 2728.2 2375.9
DeepCert ind68.target6.eps0.2 UNSAT 4.1 325.0 UNSAT 722.2 607.0
DeepCert ind68.target6.eps0.4 UNSAT 4.4 303.8 UNSAT 679.7 949.3
DeepCert ind68.target6.eps0.6 UNSAT 14.4 413.6 UNSAT 2760.5 1776.7
DeepCert ind69.target6.eps0.2 UNSAT 4.9 308.4 UNSAT 583.5 604.9
DeepCert ind69.target6.eps0.4 UNSAT 12.6 427.8 UNSAT 657.7 948.4
DeepCert ind69.target6.eps0.6 UNSAT 15.7 435.7 UNSAT 3417.4 2102.0
DeepCert ind7.target6.eps0.2 UNSAT 9.5 400.7 UNSAT 394.2 781.0
DeepCert ind7.target6.eps0.4 UNSAT 26.4 558.2 UNSAT 1031.4 1734.9
DeepCert ind7.target6.eps0.6 SAT 26.6 567.1 SAT 1410.3 2172.1
DeepCert ind70.target6.eps0.2 UNSAT 5.1 310.3 UNSAT 721.9 604.4
DeepCert ind70.target6.eps0.4 UNSAT 11.9 428.5 UNSAT 756.1 787.1
DeepCert ind70.target6.eps0.6 UNSAT 17.1 437.9 UNSAT 2683.7 1617.1
DeepCert ind71.target6.eps0.2 UNSAT 5.3 320.0 UNSAT 731.5 607.8
DeepCert ind71.target6.eps0.4 UNSAT 5.0 330.5 UNSAT 1426.9 1421.8
DeepCert ind71.target6.eps0.6 UNSAT 17.1 453.4 UNSAT 744.2 941.4
DeepCert ind72.target6.eps0.2 UNSAT 5.3 330.3 UNSAT 536.4 604.2
DeepCert ind72.target6.eps0.4 UNSAT 5.8 342.9 UNSAT 1011.7 1107.8
DeepCert ind72.target6.eps0.6 UNSAT 16.3 469.0 UNSAT 934.7 944.2
DeepCert ind73.target6.eps0.2 UNSAT 6.3 346.6 UNSAT 683.5 605.1
DeepCert ind73.target6.eps0.4 UNSAT 6.1 359.2 UNSAT 826.2 942.2
DeepCert ind73.target6.eps0.6 UNSAT 17.9 494.0 UNSAT 1616.3 1736.5
DeepCert ind74.target6.eps0.2 UNSAT 6.4 351.5 UNSAT 722.7 606.0
DeepCert ind74.target6.eps0.4 UNSAT 6.5 364.1 UNSAT 433.7 781.5
DeepCert ind74.target6.eps0.6 UNSAT 19.6 500.8 UNSAT 1150.7 1734.0
DeepCert ind75.target6.eps0.2 UNSAT 5.8 342.0 UNSAT 802.0 943.3
DeepCert ind75.target6.eps0.4 UNSAT 5.9 353.1 UNSAT 423.9 601.3
DeepCert ind75.target6.eps0.6 UNSAT 17.7 485.2 UNSAT 482.7 780.5
DeepCert ind76.target6.eps0.2 UNSAT 5.8 339.9 UNSAT 705.4 604.7
DeepCert ind76.target6.eps0.4 UNSAT 6.3 349.8 UNSAT 1254.2 941.2
DeepCert ind76.target6.eps0.6 UNSAT 15.8 481.2 UNSAT 1929.3 1617.7
DeepCert ind77.target6.eps0.2 UNSAT 5.8 342.9 UNSAT 638.5 604.6
DeepCert ind77.target6.eps0.4 UNSAT 5.7 351.0 UNSAT 787.4 780.9
DeepCert ind77.target6.eps0.6 UNSAT 5.9 359.5 UNSAT 1240.1 1452.1
DeepCert ind78.target6.eps0.2 UNSAT 7.1 340.9 UNSAT 756.0 785.4
DeepCert ind78.target6.eps0.4 UNSAT 6.7 351.8 UNSAT 556.1 603.6
DeepCert ind78.target6.eps0.6 UNSAT 18.0 483.1 UNSAT 1791.8 2415.0
DeepCert ind79.target6.eps0.2 UNSAT 7.1 353.6 UNSAT 566.5 603.6
DeepCert ind79.target6.eps0.4 UNSAT 7.7 363.0 UNSAT 2526.8 1936.0
DeepCert ind79.target6.eps0.6 UNSAT 20.8 498.7 UNSAT 2616.3 1935.7
DeepCert ind8.target6.eps0.2 UNSAT 9.5 397.3 UNSAT 510.8 941.3
DeepCert ind8.target6.eps0.4 UNSAT 34.3 552.4 UNSAT 1365.8 1495.2
DeepCert ind8.target6.eps0.6 SAT 24.7 561.8 SAT 226.7 910.8
DeepCert ind80.target6.eps0.2 UNSAT 7.9 367.0 UNSAT 527.5 604.0
DeepCert ind80.target6.eps0.4 UNSAT 7.9 376.7 UNSAT 581.7 1257.7
DeepCert ind80.target6.eps0.6 UNSAT 32.9 519.4 UNSAT 1696.3 1892.8
DeepCert ind81.target6.eps0.2 UNSAT 7.5 360.0 UNSAT 355.7 602.4
DeepCert ind81.target6.eps0.4 UNSAT 7.2 369.3 UNSAT 1711.7 2054.9
DeepCert ind81.target6.eps0.6 UNSAT 26.3 508.4 UNSAT 3017.8 2612.1
DeepCert ind82.target6.eps0.2 UNSAT 7.3 356.5 UNSAT 939.1 944.8
DeepCert ind82.target6.eps0.4 UNSAT 8.0 365.3 UNSAT 848.6 1099.6
DeepCert ind82.target6.eps0.6 UNSAT 20.4 502.4 UNSAT 1232.6 1573.7
DeepCert ind83.target6.eps0.2 UNSAT 7.0 345.7 UNSAT 619.7 604.7
DeepCert ind83.target6.eps0.4 UNSAT 7.0 354.2 UNSAT 794.1 781.3
DeepCert ind83.target6.eps0.6 UNSAT 19.2 486.6 UNSAT 1028.2 1099.1
DeepCert ind84.target6.eps0.2 UNSAT 6.3 337.1 UNSAT 541.6 603.8
DeepCert ind84.target6.eps0.4 UNSAT 6.2 348.5 UNSAT 464.5 601.9
DeepCert ind84.target6.eps0.6 UNSAT 17.6 478.1 UNSAT 1290.3 1742.3
DeepCert ind9.target6.eps0.2 UNSAT 9.6 401.3 UNSAT 335.5 605.0
DeepCert ind9.target6.eps0.4 SAT 25.3 558.1 SAT 380.7 757.5
DeepCert ind9.target6.eps0.6 SAT 26.2 567.5 SAT 761.4 3307.1
NAP cls0.id1 SAT 2.7 43.2 UNK 3601.0 1002.2
NAP cls0.id10 UNK 3601.0 57.1 UNK 3601.0 1170.2
NAP cls0.id11 SAT 1839.6 58.7 UNK 3601.0 1035.4
NAP cls0.id12 UNK 3601.0 55.7 UNK 3601.0 1035.2
NAP cls0.id13 UNK 3601.0 59.1 UNK 3601.0 1090.6
NAP cls0.id14 SAT 3.4 44.1 UNK 3601.0 952.8
NAP cls0.id15 UNK 3601.0 57.5 UNK 3601.0 1139.4
NAP cls0.id16 UNK 3601.0 59.2 UNK 3601.0 1074.2
NAP cls0.id17 SAT 179.7 56.8 UNK 3601.1 1272.1
NAP cls0.id18 UNK 3601.0 57.1 UNK 3601.1 1106.1
NAP cls0.id19 UNK 3601.0 56.6 UNK 3601.0 1132.6
NAP cls0.id2 UNK 3601.0 56.4 UNK 3601.0 1050.2
NAP cls0.id20 UNK 3601.0 58.8 UNK 3601.1 1192.9
NAP cls0.id21 SAT 2.3 42.8 UNK 3601.0 1197.1
NAP cls0.id22 UNK 3601.0 57.1 UNK 3601.0 1024.7
NAP cls0.id23 UNK 3601.0 56.5 UNK 3601.0 973.2
NAP cls0.id24 UNK 3601.0 57.4 UNK 3601.1 1241.4
NAP cls0.id25 SAT 186.1 56.5 UNK 3601.0 1087.6
NAP cls0.id26 UNK 3601.0 57.3 UNK 3601.0 984.7
NAP cls0.id27 UNK 3601.0 56.0 UNK 3601.0 1095.6
NAP cls0.id28 UNK 3601.0 57.0 UNK 3601.0 1039.9
NAP cls0.id29 UNK 3601.0 56.7 UNK 3601.0 1034.3
NAP cls0.id3 SAT 1.7 43.2 UNK 3601.0 1219.9
NAP cls0.id30 SAT 209.4 58.0 UNK 3601.0 846.8
NAP cls0.id31 SAT 1583.9 57.1 UNK 3601.0 1058.1
NAP cls0.id32 UNK 3601.0 57.2 UNK 3601.1 1208.1
NAP cls0.id33 SAT 57.8 55.6 UNK 3601.0 753.6
NAP cls0.id34 SAT 70.1 55.7 UNK 3601.0 970.2
NAP cls0.id35 UNK 3601.0 58.3 UNK 3601.0 1094.9
NAP cls0.id36 UNK 3601.0 57.4 UNK 3601.1 1128.7
NAP cls0.id37 UNK 3601.0 55.8 UNK 3601.0 1008.3
NAP cls0.id38 UNK 3601.0 56.7 UNK 3601.0 988.0
NAP cls0.id39 SAT 3.8 47.4 UNK 3601.1 1298.8
NAP cls0.id4 SAT 2.9 43.5 UNK 3601.0 1126.1
NAP cls0.id40 UNK 3601.0 56.5 UNK 3601.0 1011.1
NAP cls0.id41 UNK 3601.0 56.4 UNK 3601.0 1036.8
NAP cls0.id42 SAT 120.8 55.8 UNK 3601.1 1413.4
NAP cls0.id43 UNK 3601.0 56.8 UNK 3601.1 1215.2
NAP cls0.id44 SAT 1415.6 56.9 UNK 3601.1 1372.7
NAP cls0.id45 UNK 3601.0 57.1 UNK 3601.1 1435.9
NAP cls0.id46 UNK 3601.0 57.4 UNK 3601.0 965.0
NAP cls0.id47 UNK 3601.0 56.7 UNK 3601.1 1324.0
NAP cls0.id48 UNK 3601.0 55.4 UNK 3601.0 1104.5
NAP cls0.id49 UNK 3601.0 59.2 UNK 3601.0 928.7
NAP cls0.id5 SAT 2.0 43.8 UNK 3601.1 1152.9
NAP cls0.id50 SAT 403.0 55.9 UNK 3601.1 1138.8
NAP cls0.id51 UNK 3601.0 57.2 UNK 3601.0 998.7
NAP cls0.id52 UNK 3601.0 57.0 UNK 3601.1 1334.7
NAP cls0.id53 UNK 3601.0 59.6 UNK 3601.0 1115.1
NAP cls0.id54 UNK 3601.0 56.2 UNK 3601.1 1257.3
NAP cls0.id55 UNK 3601.0 59.8 UNK 3601.0 846.1
NAP cls0.id56 UNK 3601.0 57.1 UNK 3601.1 1247.0
NAP cls0.id57 UNK 3601.0 57.2 UNK 3601.0 1078.0
NAP cls0.id58 UNK 3601.0 58.9 UNK 3601.0 1090.7
NAP cls0.id59 UNK 3601.0 57.7 UNK 3601.0 947.0
NAP cls0.id6 UNK 3601.0 56.4 UNK 3601.0 1051.6
NAP cls0.id60 UNK 3601.0 57.1 UNK 3601.0 970.5
NAP cls0.id61 UNK 3601.0 55.5 UNK 3601.0 1106.2
NAP cls0.id62 UNK 3601.0 58.0 UNK 3601.0 912.8
NAP cls0.id63 UNK 3601.0 56.5 UNK 3601.0 1073.0
NAP cls0.id64 UNK 3601.0 58.7 UNK 3601.0 937.1
NAP cls0.id65 UNK 3601.0 57.9 UNK 3601.0 726.5
NAP cls0.id66 SAT 584.2 57.6 UNK 3601.0 1064.6
NAP cls0.id67 UNK 3601.0 55.8 UNK 3601.1 1391.0
NAP cls0.id68 UNK 3601.0 57.0 UNK 3601.0 956.8
NAP cls0.id69 UNK 3601.0 56.2 UNK 3601.1 1249.9
NAP cls0.id7 SAT 1.6 43.8 UNK 3601.0 1106.7
NAP cls0.id70 UNK 3601.0 56.8 UNK 3601.1 1110.0
NAP cls0.id71 UNK 3601.0 56.2 UNK 3601.0 1027.2
NAP cls0.id72 UNK 3601.0 56.9 UNK 3601.1 1166.1
NAP cls0.id73 SAT 1296.8 56.2 UNK 3601.0 1160.1
NAP cls0.id74 UNK 3601.0 55.4 UNK 3601.1 962.4
NAP cls0.id75 UNK 3601.0 55.9 UNK 3601.0 1137.9
NAP cls0.id76 UNK 3601.0 55.4 UNK 3601.1 1317.1
NAP cls0.id8 UNK 3601.0 57.4 UNK 3601.0 1043.1
NAP cls0.id9 SAT 2.3 43.1 UNK 3601.0 1193.6
NAP cls1.id77 UNK 3601.0 59.9 UNK 3601.1 1149.2
NAP cls1.id78 UNK 3601.0 59.4 UNK 3601.0 1102.7
NAP cls1.id79 SAT 5.2 50.7 UNK 3601.1 1157.8
NAP cls1.id80 UNK 3601.0 62.0 UNK 3601.0 941.2
NAP cls1.id81 UNK 3601.0 57.3 UNK 3601.0 981.8
NAP cls1.id82 UNK 3601.0 62.1 UNK 3601.1 1297.8
NAP cls1.id83 UNK 3601.0 58.7 UNK 3601.0 1002.1
NAP cls1.id84 SAT 1077.2 64.1 UNK 3601.0 1093.0
NAP cls1.id85 UNK 3601.0 61.2 UNK 3601.0 1135.2
NAP cls1.id86 SAT 737.2 61.5 UNK 3601.0 837.2
NAP cls1.id87 UNK 3601.0 58.3 UNK 3601.1 1222.5
NAP cls1.id88 UNK 3601.0 57.2 UNK 3601.0 1068.2
NAP cls2.id100 UNK 3601.0 57.0 UNK 3601.1 1246.2
NAP cls2.id101 UNK 3601.0 59.1 UNK 3601.0 1035.8
NAP cls2.id102 UNK 3601.0 57.6 UNK 3601.0 1107.1
NAP cls2.id103 UNK 3601.0 57.2 UNK 3601.0 1088.9
NAP cls2.id104 UNK 3601.0 56.4 UNK 3601.1 1166.8
NAP cls2.id105 UNK 3601.0 55.5 UNK 3601.0 1220.3
NAP cls2.id89 UNK 3601.0 58.3 UNK 3601.0 1030.3
NAP cls2.id90 UNK 3601.0 59.1 UNK 3601.1 1142.3
NAP cls2.id91 UNK 3601.0 55.6 UNK 3601.1 1096.6
NAP cls2.id92 UNK 3601.0 58.4 UNK 3601.1 1205.4
NAP cls2.id93 UNK 3601.0 60.4 UNK 3601.0 1057.7
NAP cls2.id94 UNK 3601.0 58.4 UNK 3601.0 634.0
NAP cls2.id95 UNK 3601.0 56.8 UNK 3601.0 881.4
NAP cls2.id96 UNK 3601.0 56.5 UNK 3601.0 1049.3
NAP cls2.id97 UNK 3601.0 56.5 UNK 3601.1 1166.9
NAP cls2.id98 UNK 3601.0 62.1 UNK 3601.0 912.4
NAP cls2.id99 UNK 3601.0 58.5 UNK 3601.1 1257.3
NAP cls3.id106 UNK 3601.0 57.6 UNK 3601.0 899.9
NAP cls3.id107 SAT 371.1 58.8 UNK 3601.1 1267.5
NAP cls3.id108 UNK 3601.0 63.3 UNK 3601.0 1093.3
NAP cls3.id109 UNK 3601.0 57.1 UNK 3601.1 1428.3
NAP cls3.id110 UNK 3601.0 62.0 UNK 3601.0 737.1
NAP cls3.id111 UNK 3601.0 59.3 UNK 3601.0 1164.7
NAP cls3.id112 SAT 3146.8 63.5 SAT 160.5 677.5
NAP cls3.id113 UNK 3601.0 57.7 UNK 3601.0 1062.7
NAP cls3.id114 UNK 3601.0 57.6 UNK 3601.1 1432.3
NAP cls3.id115 SAT 2.6 44.0 UNK 3601.1 1046.5
NAP cls3.id116 SAT 3.6 44.0 UNK 3601.0 1010.6
NAP cls3.id117 SAT 300.0 61.6 UNK 3601.0 931.4
NAP cls3.id118 SAT 13.2 53.9 UNK 3601.0 926.1
NAP cls3.id119 SAT 1019.9 61.1 UNK 3601.1 1147.0
NAP cls3.id120 UNK 3601.0 59.4 SAT 3323.4 1127.4
NAP cls3.id121 UNK 3601.0 57.4 UNK 3601.0 802.9
NAP cls3.id122 UNK 3601.0 61.2 UNK 3601.0 789.5
NAP cls3.id123 SAT 518.4 59.1 UNK 3601.0 1127.5
NAP cls3.id124 UNK 3601.0 58.0 UNK 3601.1 1268.4
NAP cls3.id125 SAT 2.6 43.7 UNK 3601.1 1013.2
NAP cls3.id126 SAT 562.3 61.2 UNK 3601.1 1405.9
NAP cls3.id127 UNK 3601.0 57.1 UNK 3601.1 1350.5
NAP cls3.id128 UNK 3601.0 60.9 UNK 3601.1 1251.3
NAP cls3.id129 UNK 3601.0 59.1 UNK 3601.0 1064.2
NAP cls3.id130 UNK 3601.0 57.4 UNK 3601.0 886.6
NAP cls3.id131 UNK 3601.0 61.4 UNK 3601.0 1146.7
NAP cls3.id132 SAT 2.8 44.3 SAT 2882.8 955.8
NAP cls3.id133 UNK 3601.0 62.7 UNK 3601.0 949.1
NAP cls3.id134 UNK 3601.0 59.6 UNK 3601.1 1016.9
NAP cls3.id135 UNK 3601.0 59.2 UNK 3601.0 1067.0
NAP cls3.id136 UNK 3601.0 59.6 UNK 3601.0 928.2
NAP cls3.id137 UNK 3601.0 58.6 UNK 3601.0 1060.5
NAP cls3.id138 UNK 3601.0 56.6 UNK 3601.1 1055.6
NAP cls3.id139 UNK 3601.0 57.3 UNK 3601.0 940.7
NAP cls3.id140 UNK 3601.0 59.3 SAT 2539.1 1029.8
NAP cls3.id141 UNK 3601.0 60.7 UNK 3601.0 1010.3
NAP cls3.id142 UNK 3601.0 59.7 UNK 3601.1 1276.4
NAP cls3.id143 UNK 3601.0 57.2 UNK 3601.1 1107.5
NAP cls3.id144 UNK 3601.0 57.3 UNK 3601.0 980.8
NAP cls3.id145 UNK 3601.0 57.9 SAT 81.3 905.3
NAP cls3.id146 UNK 3601.0 61.9 SAT 842.4 933.9
NAP cls3.id147 UNK 3601.0 57.2 UNK 3601.1 1170.6
NAP cls3.id148 UNK 3601.0 62.0 UNK 3601.0 1140.8
NAP cls3.id149 UNK 3601.0 57.0 UNK 3601.0 978.0
NAP cls3.id150 UNK 3601.0 57.2 UNK 3601.1 1595.4
NAP cls3.id151 UNK 3601.0 57.7 UNK 3601.0 896.0
NAP cls3.id152 UNK 3601.0 57.3 UNK 3601.0 1099.5
NAP cls3.id153 UNK 3601.0 62.1 UNK 3601.0 995.1
NAP cls3.id154 UNK 3601.0 57.3 UNK 3601.0 1167.0
NAP cls3.id155 UNK 3601.0 57.3 UNK 3601.0 1098.2
NAP cls3.id156 UNK 3601.0 55.1 UNK 3601.0 1009.3
NAP cls3.id157 UNK 3601.0 57.0 UNK 3601.0 1106.1
NAP cls3.id158 UNK 3601.0 57.1 SAT 49.0 625.2
NAP cls3.id159 UNK 3601.0 58.1 UNK 3601.0 1099.7
NAP cls3.id160 UNK 3601.0 61.4 UNK 3601.0 963.7
NAP cls3.id161 UNK 3601.0 59.3 UNK 3601.1 950.2
NAP cls3.id162 UNK 3601.0 57.0 UNK 3601.0 1014.4
NAP cls3.id163 UNK 3601.0 61.0 UNK 3601.0 1058.5
NAP cls3.id164 UNK 3601.0 56.8 UNK 3601.0 1028.0
NAP cls3.id165 UNK 3601.0 57.4 UNK 3601.1 1209.0
NAP cls3.id166 SAT 3.7 44.2 UNK 3601.1 1132.5
NAP cls3.id167 UNK 3601.0 57.0 UNK 3601.0 962.5
NAP cls3.id168 SAT 1130.6 59.8 UNK 3601.1 1217.6
NAP cls3.id169 UNK 3601.0 62.1 UNK 3601.1 1200.4
NAP cls3.id170 UNK 3601.0 59.4 UNK 3601.0 820.0
NAP cls3.id171 UNK 3601.0 61.2 UNK 3601.1 1468.8
NAP cls3.id172 UNK 3601.0 57.0 UNK 3601.0 993.7
NAP cls4.id173 UNK 3601.0 56.8 UNK 3601.0 907.6
NAP cls4.id174 UNK 3601.0 58.7 UNK 3601.0 975.6
NAP cls4.id175 UNK 3601.0 59.5 UNK 3601.1 1343.9
NAP cls4.id176 UNK 3601.0 58.9 UNK 3601.0 1095.5
NAP cls4.id177 UNK 3601.0 56.9 UNK 3601.1 1105.3
NAP cls4.id178 UNK 3601.0 59.1 UNK 3601.0 982.6
NAP cls4.id179 UNK 3601.0 56.8 UNK 3601.0 1010.6
NAP cls4.id180 SAT 1157.6 63.1 UNK 3601.1 1028.0
NAP cls4.id181 UNK 3601.0 56.9 UNK 3601.1 1180.4
NAP cls4.id182 UNK 3601.0 57.7 SAT 171.3 530.2
NAP cls4.id183 UNK 3601.0 55.0 UNK 3601.0 890.4
NAP cls4.id184 UNK 3601.0 61.1 UNK 3601.0 968.7
NAP cls4.id185 UNK 3601.0 60.8 UNK 3601.1 1083.6
NAP cls4.id186 UNK 3601.0 57.3 UNK 3601.0 1089.6
NAP cls4.id187 UNK 3601.0 56.9 UNK 3601.1 1143.1
NAP cls4.id188 UNK 3601.0 58.2 UNK 3601.1 1324.8
NAP cls4.id189 UNK 3601.0 57.5 UNK 3601.1 1257.3
NAP cls4.id190 UNK 3601.0 56.5 UNK 3601.0 1255.5
NAP cls4.id191 UNK 3601.0 57.8 UNK 3601.1 1081.5
NAP cls4.id192 UNK 3601.0 56.1 UNK 3601.0 986.5
NAP cls4.id193 UNK 3601.0 56.8 UNK 3601.1 1240.3
NAP cls4.id194 UNK 3601.0 60.9 UNK 3601.0 824.6
NAP cls4.id195 SAT 2.3 44.5 UNK 3601.1 1105.0
NAP cls4.id196 UNK 3601.0 59.2 UNK 3601.0 1081.9
NAP cls4.id197 UNK 3601.0 56.5 UNK 3601.1 1350.3
NAP cls4.id198 SAT 2631.4 61.6 UNK 3601.1 1241.2
NAP cls4.id199 UNK 3601.0 57.0 UNK 3601.0 793.0
NAP cls4.id200 UNK 3601.0 57.6 UNK 3601.0 1015.7
NAP cls4.id201 UNK 3601.0 59.1 UNK 3601.0 997.9
NAP cls4.id202 UNK 3601.0 57.0 UNK 3601.0 812.3
NAP cls4.id203 UNK 3601.0 60.1 UNK 3601.1 1084.9
NAP cls4.id204 UNK 3601.0 56.4 UNK 3601.1 1514.7
NAP cls4.id205 UNK 3601.0 60.5 UNK 3601.0 967.2
NAP cls4.id206 UNK 3601.0 56.8 UNK 3601.0 949.8
NAP cls4.id207 UNK 3601.0 56.7 UNK 3601.1 1294.3
NAP cls4.id208 UNK 3601.0 56.7 UNK 3601.1 961.4
NAP cls4.id209 UNK 3601.0 56.9 UNK 3601.1 1251.7
NAP cls4.id210 UNK 3601.0 59.1 UNK 3601.0 774.3
NAP cls4.id211 UNK 3601.0 57.1 UNK 3601.0 992.3
NAP cls4.id212 UNK 3601.0 57.0 UNK 3601.1 1356.3
NAP cls4.id213 UNK 3601.0 58.9 UNK 3601.0 1095.2
NAP cls4.id214 UNK 3601.0 56.9 UNK 3601.1 1091.5
NAP cls4.id215 UNK 3601.0 58.4 UNK 3601.1 1154.9
NAP cls4.id216 UNK 3601.0 57.2 UNK 3601.1 1044.2
NAP cls4.id217 UNK 3601.0 56.2 UNK 3601.0 887.3
NAP cls4.id218 UNK 3601.0 58.2 UNK 3601.1 1367.5
NAP cls4.id219 UNK 3601.0 58.7 UNK 3601.1 1085.1
NAP cls4.id220 UNK 3601.0 56.3 UNK 3601.1 1167.6
NAP cls4.id221 UNK 3601.0 57.1 UNK 3601.0 1051.8
NAP cls4.id222 UNK 3601.0 58.9 UNK 3601.1 1407.6
NAP cls4.id223 SAT 267.4 61.3 UNK 3601.0 1072.7
NAP cls4.id224 UNK 3601.0 57.7 UNK 3601.1 1041.2
NAP cls4.id225 UNK 3601.0 56.0 UNK 3601.1 1136.6
NAP cls4.id226 UNK 3601.0 54.0 UNK 3601.1 1185.4
NAP cls4.id227 UNK 3601.0 58.8 UNK 3601.1 1159.9
NAP cls4.id228 UNK 3601.0 56.3 UNK 3601.1 1199.0
NAP cls4.id229 UNK 3601.0 56.7 UNK 3601.1 1493.2
NAP cls4.id230 UNK 3601.0 56.4 UNK 3601.0 937.2
NAP cls4.id231 UNK 3601.0 57.1 UNK 3601.0 822.6
NAP cls4.id232 UNK 3601.0 57.3 UNK 3601.1 1199.7
NAP cls4.id233 UNK 3601.0 56.9 UNK 3601.0 973.0
NAP cls4.id234 UNK 3601.0 57.8 SAT 15.6 265.8
NAP cls4.id235 UNK 3601.0 56.0 UNK 3601.1 1155.8
VeriX ind0 OK 571.7 778.7 OK 1483.1 1048.9
VeriX ind1 OK 568.9 790.1 OK 1504.1 1045.5
VeriX ind10 OK 670.7 528.1 OK 1665.6 501.9
VeriX ind100 OK 1368.8 528.1 UNK 3601.0 502.0
VeriX ind101 OK 3230.5 528.4 UNK 3601.0 503.0
VeriX ind102 OK 630.0 529.6 OK 1683.7 502.6
VeriX ind103 OK 989.4 527.7 UNK 3601.0 501.8
VeriX ind11 OK 1456.7 529.3 UNK 3601.0 502.2
VeriX ind12 OK 775.5 528.3 OK 2642.2 1048.7
VeriX ind13 OK 1777.8 966.2 UNK 3601.0 502.0
VeriX ind14 OK 674.8 1101.1 OK 1556.6 503.3
VeriX ind15 OK 635.6 527.4 OK 1791.4 502.0
VeriX ind16 OK 1210.9 527.6 OK 2818.4 501.8
VeriX ind17 OK 543.7 527.9 OK 1449.7 502.3
VeriX ind19 OK 2235.1 527.7 UNK 3601.0 502.1
VeriX ind2 OK 1174.1 581.2 OK 3257.5 1045.6
VeriX ind20 OK 667.8 528.5 OK 2091.4 1050.2
VeriX ind21 OK 762.4 528.0 OK 2303.0 502.5
VeriX ind22 OK 2646.4 528.5 UNK 3601.0 503.4
VeriX ind23 OK 666.4 528.2 OK 1715.9 502.1
VeriX ind24 OK 2386.4 528.4 UNK 3601.0 501.9
VeriX ind25 OK 1662.5 528.8 OK 2197.4 502.5
VeriX ind26 OK 759.5 528.1 OK 2392.8 502.4
VeriX ind27 OK 982.9 528.0 UNK 3601.0 501.8
VeriX ind28 OK 636.4 528.1 OK 1961.3 501.8
VeriX ind29 OK 633.3 528.4 OK 1484.9 501.8
VeriX ind3 OK 936.6 966.8 OK 2748.7 1045.5
VeriX ind30 OK 667.1 528.3 OK 1703.9 502.7
VeriX ind31 OK 1168.0 528.4 OK 1955.8 501.9
VeriX ind32 OK 603.8 528.7 OK 1627.3 501.8
VeriX ind34 OK 591.5 528.4 OK 1409.0 502.1
VeriX ind35 OK 816.2 528.7 OK 1722.8 501.6
VeriX ind36 OK 848.3 528.6 OK 1508.8 501.2
VeriX ind37 OK 1212.0 528.5 OK 1942.0 502.3
VeriX ind38 OK 429.9 528.2 OK 1100.8 502.4
VeriX ind39 OK 1108.5 528.2 OK 2102.3 502.3
VeriX ind4 OK 1615.9 1098.9 UNK 3601.0 502.1
VeriX ind40 UNK 3601.0 529.6 UNK 3601.0 501.7
VeriX ind41 OK 1381.1 528.1 OK 1641.0 502.0
VeriX ind42 OK 737.1 529.1 OK 2096.1 502.0
VeriX ind43 OK 1199.7 528.4 OK 3025.6 502.0
VeriX ind44 OK 530.9 528.0 OK 1329.1 501.7
VeriX ind45 OK 703.2 529.5 OK 1860.3 502.2
VeriX ind46 OK 537.4 528.8 OK 1274.5 501.6
VeriX ind47 OK 633.7 528.0 OK 1715.8 502.6
VeriX ind48 OK 632.7 528.7 OK 1684.1 502.9
VeriX ind49 OK 3228.2 528.2 UNK 3601.0 501.6
VeriX ind5 OK 1135.8 1099.1 OK 3101.4 1047.2
VeriX ind50 OK 1210.4 528.4 OK 2965.5 501.6
VeriX ind51 OK 681.8 528.8 OK 1804.2 502.2
VeriX ind52 OK 730.1 528.2 OK 1957.1 502.0
VeriX ind53 OK 1615.6 529.1 UNK 3601.0 501.6
VeriX ind54 OK 651.7 528.2 OK 1567.8 502.6
VeriX ind55 OK 664.8 528.4 OK 1829.7 501.7
VeriX ind56 OK 1014.9 528.1 UNK 3600.0 501.5
VeriX ind57 OK 1403.4 528.8 OK 2715.4 501.6
VeriX ind58 OK 718.6 528.9 OK 2200.0 502.4
VeriX ind59 OK 3330.6 528.8 UNK 3601.0 501.6
VeriX ind6 OK 655.4 528.6 OK 2558.8 1041.0
VeriX ind60 OK 579.6 528.0 OK 1638.1 501.5
VeriX ind61 OK 431.9 527.4 OK 1103.4 503.1
VeriX ind62 OK 483.9 543.7 OK 1212.7 502.2
VeriX ind64 OK 581.6 528.4 OK 1476.5 501.7
VeriX ind65 OK 628.2 528.1 OK 2251.1 502.0
VeriX ind66 OK 604.7 528.3 OK 2301.6 501.0
VeriX ind67 OK 1415.0 528.9 UNK 3601.0 501.8
VeriX ind68 OK 665.7 528.9 OK 1825.2 502.1
VeriX ind69 OK 614.5 528.4 OK 1700.0 502.3
VeriX ind7 OK 656.8 1092.9 OK 1854.0 509.2
VeriX ind70 OK 1254.9 528.9 OK 1947.7 502.8
VeriX ind71 OK 728.4 528.0 OK 1972.6 502.4
VeriX ind72 OK 652.7 528.1 OK 1701.5 501.1
VeriX ind73 OK 871.9 528.5 OK 2224.3 502.1
VeriX ind74 OK 1096.8 528.2 OK 1902.1 501.9
VeriX ind75 OK 826.6 527.8 OK 1841.3 501.3
VeriX ind76 OK 647.8 528.0 OK 1898.4 501.2
VeriX ind77 OK 3181.5 529.1 UNK 3601.0 502.6
VeriX ind78 OK 631.6 545.2 OK 1777.7 501.9
VeriX ind79 OK 595.8 528.0 OK 1527.9 502.3
VeriX ind80 OK 618.7 544.5 OK 1633.8 502.4
VeriX ind81 OK 762.3 528.3 OK 2279.7 502.1
VeriX ind82 OK 633.9 528.1 OK 1741.3 502.1
VeriX ind83 OK 693.7 527.9 OK 1831.7 501.8
VeriX ind84 OK 1008.4 528.7 OK 3061.4 502.0
VeriX ind85 OK 664.7 528.0 OK 1853.2 502.1
VeriX ind86 OK 889.0 529.4 OK 1896.5 501.7
VeriX ind87 OK 488.7 544.5 OK 1265.4 502.9
VeriX ind88 OK 1373.3 527.7 OK 2453.0 501.4
VeriX ind89 OK 1076.0 527.4 OK 2290.9 503.0
VeriX ind9 OK 796.1 532.4 OK 2213.8 502.1
VeriX ind90 OK 649.4 529.3 OK 1842.2 502.9
VeriX ind91 OK 754.0 528.4 OK 2204.0 501.3
VeriX ind92 OK 589.6 528.5 OK 1653.2 501.5
VeriX ind93 OK 617.3 527.3 OK 1608.4 502.3
VeriX ind94 OK 692.8 528.5 OK 1637.0 502.8
VeriX ind95 OK 614.4 529.1 OK 1863.5 503.2
VeriX ind96 OK 1197.3 528.0 OK 1689.1 502.4
VeriX ind97 OK 373.8 545.5 OK 979.6 501.7
VeriX ind98 OK 600.4 528.7 OK 1500.1 501.9
VeriX ind99 OK 1290.3 527.9 OK 3130.3 501.8